### v1

# DualContrast: Unsupervised Disentangling of Content and Transformations with Implicit Parameterization

Abstract
Unsupervised disentanglement of content and transformation has recently drawn much research, given their efficacy in solving downstream unsupervised tasks like clustering, alignment, and shape analysis. This problem is particularly important for analyzing shape-focused real-world scientific image datasets, given their significant relevance to downstream tasks. The existing works address the problem by explicitly parameterizing the transformation factors, significantly reducing their expressiveness. Moreover, they are not applicable in cases where transformations can not be readily parametrized. An alternative to such explicit approaches is self-supervised methods with data augmentation, which implicitly disentangles transformations and content. We demonstrate that the existing self-supervised methods with data augmentation result in the poor disentanglement of content and transformations in real-world scenarios. Therefore, we developed a novel self-supervised method, DualContrast, specifically for unsupervised disentanglement of content and transformations in shape-focused image datasets. Our extensive experiments showcase the superiority of DualContrast over existing self-supervised and explicit parameterization approaches. We leveraged DualContrast to disentangle protein identities and protein conformations in cellular 3D protein images. Moreover, we also disentangled transformations in MNIST, viewpoint in the Linemod Object dataset, and human movement deformation in the Starmen dataset as transformations using DualContrast.
Conceptualization of content-transformation disentanglement (a) and comparison between existing explicit (b) and our implicit (c) unsupervised methods. Both of these methods optimize content and transformation latent spaces while reconstructing $\mathbf{x}^{\prime}$ from $\mathbf{x}$. However, explicit methods use the transformation space to infer a fixed parameter set $\gamma$ and use it to transform the feature map/image $\mathbf{x}_{c}$ generated by conditioning the decoder on content space. On the contrary, our implicit method conditions the decoder on both content and transformation to generate image $\mathbf{x}^{\prime}$ and does not restrict the transformation space to a fixed set of parameters.
1 Introduction
Figure 1: Our proposed contrastive learning-based unsupervised content-transformation disentanglement pipeline. (a) The variational inference of content and transformation factors with $L_{\text{VAE}}$. (b,c) The contrastive strategy to create positive and negative pairs w.r.t. content (b) and transformation factors (c), respectively. DualContrast combines the three parts and is trained end-to-end.
In many real-world image datasets, particularly in scientific imaging domains, the object shapes being visualized may undergo multiple transformations [1, 2]. These shape-focused image datasets can thus be regarded as samples generated from two independent factors, one representing the semantic attribute termed content and the other representing the transformations. Content is regarded as the form of an object being visualized in the image [1] that is unaltered after applying any nuisance transformation. On the other hand, transformation dictates the specific state or realization of that form of the object. Taking proteins inside the cells as an example, changing the transformation factor changes the conformation of the protein where the protein identity (i.e., protein composition) does not change (Figure DualContrast: Unsupervised Disentangling of Content and Transformations with Implicit Parameterization). On the other hand, changing the content factor is analogous to changing the proteins from one identity to another through compositional change. Such a phenomenon holds for many real-world shape-focused datasets. Disentangling content and transformation factors of images can significantly facilitate several downstream tasks, including clustering [3], alignment [4, 2], zero-shot transfer [5], bias removal [6, 7], cross-domain retrieval [8, 9], shape analysis [10, 11, 4, 2, 3], etc. [11].
A number of works have been done to directly or indirectly solve the content-transformation disentangling problem [12, 8, 10, 1, 4, 2], whereas many of them use annotated factors for training. However, such annotation is hard to obtain, and thus, unsupervised disentanglement methods are desired. Only a few works [1, 2, 4] have addressed the problem in a completely unsupervised setting. Among them, VITAE [1] performed explicit parameterization of transformation factors as diffeomorphic transformations and achieved noteworthy disentanglement of content and transformation in several image datasets. SpatialVAE [2] and Harmony [4] focus on representing known transformation types as transformation factors. SpatialVAE [2] performed disentanglement of two-dimensional translation and rotation, whereas Harmony [4] demonstrated disentanglement of a broader set of transformations that can be explicitly parameterized.
Though these unsupervised explicit parameterization methods have demonstrated several successes, they have several major limitations in general: (1) These method uses the transformation factor to generate only a few parameterized transformations. In reality, many transformations do not have a well-defined parameterized form,   e.g., protein conformation changes, viewpoint change (LineMod), human deformation during movement (Starmen), etc. These explicit methods can not disentangle such transformation by design. (2) They use Spatial Transformer Networks (STN) [13] for inferring the transformation factor, which requires a continuous parameterization of the transformation to be disentangled. The continuity of many transformations (e.g., SO(3) rotation) in neural networks is often a concern [14]. Moreover, (3) these methods often infer spurious contents during disentanglement in real-world datasets(Fig. 2). Due to these limitations, an unsupervised content-transformation disentangling method without explicit transformation representation is much desired.
Consequently, in this work, we focused on the more generalized setting of unsupervised content-transformation disentanglement without any explicit parameterization of the transformation factor. Given the theoretical study by Von et al. [15] showing the implicit disentanglement ability of standard Contrastive Learning (CL) methods, e.g., SimCLR [16], to disentangle content from style, a plausible solution could be using SimCLR with geometric data augmentation for our implicit content-transformation disentanglement task. However, these popular CL methods [16] have rarely achieved disentanglement in practice [6], and we also found that SimCLR with geometric data augmentation performs poorly in our shape-focused image datasets of interest (Table 1).
To this end, we develop a novel contrastive learning method for implicit content-transformation disentanglement, termed “DualContrast”. Similar to previous content-transformation disentangling methods [17, 2, 1, 4], it is a deep generative model based on a variational autoencoder (VAE) architecture. However, unlike previous methods, it does not require labels or explicit parameterization of any latent factor. Instead, it uses a novel contrastive strategy that creates “positive” and “negative” contrastive pairs w.r.t. both content and transformation latent factor. We create negative pairs w.r.t. transformation factor and positive pairs w.r.t. content factor by simply rotating the samples. In addition, we randomly permute the samples in the batch to create a negative pair for a batch of samples w.r.t. content latent factor. Without prior knowledge, creating positive pairs w.r.t. the transformation latent factor from the input samples is not possible. To this end, we generate samples from an identical distribution of the transformation representations in the embedding space and treat them as positive pairs for the transformation latent factor. Upon creating both “positive” and “negative” pairs w.r.t. both factors, we penalize similarity between negative pairs and distance between positive pairs for each factor. In DualContrast, we combine these contrastive losses with VAE reconstruction loss and train the model in an end-to-end manner.
With such a strategy, by just performing simple rotation to create contrastive pairs, DualContrast disentangles a much wider range of more complicated transformations, e.g., human deformation in starmen, viewpoint in Linemod, the writing style in MNIST, etc. With DualContrast, we, for the first time, solved the scientific problem of disentangling protein identities from protein conformations in simulated 3D protein images (Figure 6) inside the cell, known as cellular cryo-ET subtomograms. DualContrast could successfully disentangle the protein identities as content and protein conformations as transformation factors in simulated subtomograms, which is not achievable with previous methods. Overall, our qualitative and quantitative experimental results show the efficacy of DualContrast in isolating content from transformation compared to existing relevant methods.
A summary of our contributions is as follows:
  • We work on a challenging setting of learning unsupervised disentangled representations of content and transformation in shape-focused images by removing the requirement of explicit parameterization.
  • We develop a novel contrastive generative modeling strategy that creates positive-negative pairs for both content and transformation latent factors. We introduced a novel way of designing contrastive pairs for transformation factors.
  • We show that DualContrast is highly effective in disentangling a wide range of transformations from the content in various shape-focused image datasets by only using simple rotation for creating contrastive pairs.
  • Leveraging DualContrast, we, for the first time, disentangled protein identities (i.e.composition) from protein conformations in simulated 3D cellular cryo-ET subtomograms, as a proof of principle.
2 Related Works
Disentangled Representation Learning: Our method can be regarded as a disentangled representation learning method that disentangles generative factors of variation as content-specific and transformation-specific factors. There exist many generic disentangled representation learning approaches [18, 17, 19, 20, 21, 22] where the factors of variations are interpreted by traversing through the learned factors and reconstructing samples. However, these methods do not perform very well in specifically disentangling content and transformation factors [1]. To this end, specialized methods that deal with such two factors have been introduced recently. Many of them [23, 12, 8] are partially or weakly supervised and use labels associated with one of the latent factors to train their models. However, these methods are not applicable when such supervision is unavailable beforehand, which is often the case for many real-life datasets. There are a few [1, 4, 2, 15] approaches that deal with the unsupervised disentanglement of content and transformation. However, these methods perform explicit parameterization of the transformation factor as affine or some specific parameterized transformation. In contrast, our method does not impose any explicit parameterization on the transformation latent factor and does not face the issues of the explicit methods (Section 1).
Contrastive Learning based Disentanglement: Contrastive learning is the primary building block of our method. Contrastive Learning with data augmentation with existing approaches, e.g., InfoNCE loss-based SimCLR [16], etc., has been previously used by [15, 8] to isolate content from style. Though in several scenarios, style in shape focused images can be referred to as transformations, these works do not specifically focus on shape-focused real-world images. Moreover, our experiments demonstrate that SimCLR with rotation augmentation in a standard encoder-decoder framework leads to very poor disentanglement of content and transformation (Table 1). Moreover, [15] only creates positive pairs w.r.t content and negative pair w.r.t. style with data augmentation in SimCLR. We have observed this scheme does not work well in our scenario. Unlike these works, our approach uses a novel strategy to create “negative" and “positive" pairs for both content and transformation of latent factors without any InfoNCE loss.
Unsupervised Content-style disentanglement: Apart from [15], several methods [6, 24, 25, 26, 27, 28] exist that perform unsupervised content-style disentanglement, focusing on natural images. These methods focus on the controllable generation of realistic natural images, mostly using ImageNet pretrained models. Unlike these methods, our work primarily focuses on disentangling content and transformation in shape-focused image datasets. Moreover, we do not depend on any ImageNet pretrained models as our images of interest differ greatly from the natural images of ImageNet.
Protein Composition-Conformation Disentanglement: One of the major contributions of our work is disentangling protein identity characterized by large compositional variability as the content factor and protein states characterized by conformational variability or subtle compositional variability as transformation factor from 3D cellular cryo-electron tomography (cryo-ET) subimages or subtomograms. Previously, Harmony [4] was used to disentangle large compositional variability in cellular subtomograms. HEMNMA-3D [29] and TomoFlow [30] were used to disentangle subtle compositional or conformational variability in subtomograms using known templates. SpatialVAE [2] disentangled subtle conformational variability from 2D cryo-EM images. There exists several works on single particle cryo-EM and cryo-ET reconstruction, e.g., cryoDRGN2 [31], cryoFIRE [3], cryoAI [32], etc. that performs amortized inference of transformation ($SO(3)\times d^{2}$) and latent space representing content. However, these works mainly focus on 2D-to-3D reconstruction instead of content-transformation disentanglement. Ours is the first work to deal with both large and subtle compositional variability in two different latent spaces for 3D cryo-ET subtomograms.
3 Method
3.1 Disentangling Content and Transformation
Disentangling content and transformation refers to learning one mapping $h:\mathcal{X}\rightarrow\mathcal{C}\times\mathcal{Z}$, where $\mathcal{X}$ is an input space, $\mathcal{C}$ and $\mathcal{Z}$ are the intermediary content and transformation representation spaces respectively. Here, $h$ can be decomposed as $h_{c}:\mathcal{X}\rightarrow\mathcal{C}$ and $h_{z}:\mathcal{X}\rightarrow\mathcal{Z}$ representing the content and transformation mapping, respectively. In an unsupervised setting, distinguishing between content and transformation can be tricky. In this paper, we distinguish between content and transformation with respect to a family of transformations $\mathcal{T}:\mathcal{X}\rightarrow\mathcal{X}$ such that the following two conditions hold:
  • Condition 1: $\forall$ $T\in\mathcal{T}$, s.t., for any $\mathbf{x}\in\mathcal{X}$, $h_{c}(T(\mathbf{x}))=h_{c}(\mathbf{x})$.
  • Condition 2: $\exists$ $T\in\mathcal{T}$, s.t., for any $\mathbf{x}^{(1)},\mathbf{x}^{(2)}\in\mathcal{X}$, $h_{z}(T(\mathbf{x}^{(1)}))=h_{z}(\mathbf{x}^{(2)})$.
Condition 1 defines content space which is invariant w.r.t. $\mathcal{T}$. On the other hand, condition 2 states that there is a $T\in\mathcal{T}$ that can make the transformation of any two samples equal when applied to any one of them. Thus it defines transformation space which is informative of $\mathcal{T}$.
For content-transformation disentangling, changes in $\mathcal{C}$ must be unaffected by changes in $\mathcal{Z}$ and vice versa. In our problem setting, no labels associated with content space $C$ or transformation space $Z$ are known. The exact family of transformation $\mathcal{T}$ is also not known beforehand. Thus, we aim to simultaneously learn the family of transformations $\mathcal{T}$ and the content factors unaltered by $\mathcal{T}$.
3.2 Method Overview and Notation
Our method is built upon a variational auto-encoder (VAE) based architecture (Figure 1), where we perform inference of the content and transformation factors in the data. Consider, a sample space $\mathcal{X}$, two latent distribution spaces $\Phi_{Z}$ and $\Phi_{C}$, two latent spaces $\mathcal{Z}$ and $\mathcal{C}$. We apply an encoder network $q_{\psi}:\mathcal{X}\rightarrow\mathcal{C}\times\mathcal{Z}$ on a sample space $bx\in\mathcal{X}$ to encode content distribution $\phi_{c}\in\Phi_{c}$ and transformation distribution $\phi_{z}\in\Phi_{z}$, where $(\phi_{c},\phi_{z})=q_{\psi}(\mathbf{x})$. Latent parameters $\mathbf{c}\in\mathcal{C}$, $\mathbf{z}\in\mathcal{Z}$ are drawn from distributions $\phi_{c}$ and $\phi_{z}$ respectively. These parameters are then decoded using a decoder network $p_{\theta}:\mathcal{C}\times\mathcal{Z}\rightarrow\mathcal{X}$ to produce $x_{\text{recon}}=p_{\theta}(\mathbf{c},\mathbf{z})$.
3.3 Variational Inference of Content and Transformation Factors
Similar to a variational auto-encoder (VAE) setting, the encoder and decoder network parameters are simultaneously optimized to maximize the variational lower bound to the likelihood $p(\mathbf{x})$ called the evidence lower bound (ELBO). ELBO maximizes the log-likelihood of the data and minimizes the KL divergence between latent distributions and a prior distribution. The objective function for this can be written as:
$\begin{split}p(\mathbf{x})\geq&\mathbb{E}_{q_{\psi}(\mathbf{c},\mathbf{z}|\mathbf{x})}[\log\frac{p_{\theta}(\mathbf{x},\mathbf{c},\mathbf{z})}{q_{\psi}(\mathbf{c},\mathbf{z}|\mathbf{x})}]\\ &=\mathbb{E}_{q_{\psi}(\mathbf{c},\mathbf{z}|\mathbf{x})}\log p_{\theta}(\mathbf{x}|\mathbf{c},\mathbf{z})-KL(q_{\psi}(\mathbf{c}|\mathbf{x})||p(\mathbf{c}))-KL(q_{\psi}(\mathbf{z}|\mathbf{x})||p(\mathbf{z})).\end{split}$ (1)
The log-likelihood term is maximized using a reconstruction loss $L_{\text{recon}}$. The prior for $\mathbf{c}$ and $\mathbf{z}$, $p(\mathbf{c})$ and $p(\mathbf{z})$ respectively are assumed to be standard normal distribution. In short, we write the KL terms as $L_{\text{KL}(c)}$ and $L_{\text{KL}(z)}$. To regularize the effect of prior on the final objective separately for two latent factors, we scale the two losses with two hyperparameters $\gamma_{c}$ and $\gamma_{z}$. Our loss function for variational inference becomes:
$L_{\text{vae}}=L_{\text{rec}}+\gamma_{c}L_{\text{KL}(\mathbf{c})}+\gamma_{z}L_{\text{KL}(\mathbf{z})}.$ (2)
We observe that by deactivating the KL term associated with any one of the factors, the likelihood tends to be optimized through the other latent factor. For instance, if $\gamma_{c}$ is set as 00 and $\gamma_{z}$ has a large value, then the $L_{\text{rec}}$ is optimized through $\mathbf{c}$ only. As a consequence, $\mathbf{c}$ captures all the information in $\mathbf{x}$, whereas $\mathbf{z}$ becomes totally independent of the data. In such a scenario, $\mathbf{z}$ will be a random factor with no relation to the transformation factor of the data, which is not desirable.
3.4 Creating Contrastive Pair with respect to Content Latent Factor
We adopt a contrastive strategy for content-transformation disentanglement. We first develop a strategy to impose a structure to the content latent factor with contrastive learning. To this end, we create negative and positive pairs for each data sample w.r.t. the content factor. Hence, we randomly pick two samples $\mathbf{x}$ and $\mathbf{x}_{\text{neg}(\mathbf{c})}$, where $\mathbf{x},\mathbf{x}_{\text{neg}(\mathbf{c})}\in\mathcal{X}$ and regard them as negative pairs of each other w.r.t the content factor. Such strategy is commonly used in traditional self-supervised learning approaches [16, 33]. Considering the high heterogeneity present in the dataset, evidenced by a large number of classes, this approach accounts for the likelihood that any two randomly selected samples will have different content. To generate positive pairs w.r.t. the content factor, we pick any sample $\mathbf{x}$ and transform it with a transformation $T\in\mathcal{T}$ to generate $\mathbf{x}_{\text{pos}(\mathbf{c})}$. We use the encoder $q_{\psi}$ to generate content factors $\mathbf{c}$, $\mathbf{c}_{\text{pos}}$, and $\mathbf{c}_{\text{neg}}$ respectively from $\mathbf{x}$, $\mathbf{x}_{\text{pos}(\mathbf{c})}$, and $\mathbf{x}_{\text{neg}(\mathbf{c})}$. We use a contrastive loss to penalize the distance between $\mathbf{c}$ and $\mathbf{c}_{\text{pos}}$ and the similarity between $\mathbf{c}$ and $\mathbf{c}_{\text{neg}}$. Our contrastive loss for content latent factor can be written as:
$L_{\text{con}(\mathbf{c})}=L_{\text{dist}}(\mathbf{c},\mathbf{c}_{\text{pos}})+L_{\text{sim}}(\mathbf{c},\mathbf{c}_{\text{neg}})$ (3)
To implement $L_{\text{dist}}$ and $L_{\text{sim}}$, we use mean absolute cosine distance and mean absolute cosine similarity, respectively.
3.5 Creating Contrastive Pair with respect to Transformation Latent Factor
Imposing contrastive loss on the content latent factor satisfies condition 1 of the content-transformation disentanglement (Section 3.1). However, the content-transformation disentanglement remains only partial since it does not satisfy condition 2. To solve this issue, we design another contrastive loss w.r.t. the transformation latent factor. Designing a contrastive loss that explicitly enforces condition 2 is not possible, given the transformation that may equalize the transformation factor of two samples can not be approximated. To this end, we design a contrastive loss that implicitly encourages condition 2. We validated the design experimentally in Section 4.
The negative pair of a sample $\mathbf{x}\in\mathcal{X}$ w.r.t. transformation latent factor is generated while creating a positive pair of it w.r.t content factor. This is because a transformation $T$ on $\mathbf{x}$ alters its transformation, but not its contents. So, we regard $\mathbf{x}_{\text{pos}(\mathbf{c})}$ as $\mathbf{x}_{\text{neg}(\mathbf{z})}$. However, creating positive pairs of samples w.r.t. transformation is very challenging as they can not be created in the input space, unlike others. To this end, we adopted a strategy of creating pairs of samples from the same distribution in the transformation factor embedding space. We randomly sample $\mathbf{z}^{(1)}$ and $\mathbf{z}^{(2)}$ from $\mathcal{N}(0,1)$ and use the decoder $p_{\theta}$ to generate samples $\mathbf{x}^{(1)}_{\text{pos}(\mathbf{z})}$ and $\mathbf{x}^{(2)}_{\text{pos}(\mathbf{z})}$ from them (Fig. 1 (c)). In the decoder, as content input, we use content factor $\mathbf{c}^{(1)}$ and $\mathbf{c}^{(2)}$ obtained by encoding two random samples $\mathbf{x}^{(1)}$ and $\mathbf{x}^{(2)}$. Consequently, $\mathbf{x}^{(1)}_{\text{pos}(\mathbf{z})}=p_{\theta}(\mathbf{c}^{(1)},\mathbf{z}^{(1)})$ and $\mathbf{x}^{(2)}_{\text{pos}(\mathbf{z})}=p_{\theta}(\mathbf{c}^{(2)},\mathbf{z}^{(2)})$ serve as the positive pairs w.r.t. transformation latent factor. We then use the encoder $q_{\psi}$ to generate $\mathbf{z}$, $\mathbf{z}_{\text{neg}}$, $\mathbf{z}^{(1)}_{\text{pos}}$, and $\mathbf{z}^{(2)}_{\text{pos}}$ from $\mathbf{x}$, $\mathbf{x}_{\text{neg}(\mathbf{z})}$, $\mathbf{x}^{(1)}_{\text{pos}(\mathbf{z})}$, and $\mathbf{x}^{(2)}_{\text{pos}(\mathbf{z})}$ respectively. We use a contrastive loss to penalize the distance between $\mathbf{z}^{(1)}_{\text{pos}}$ and $\mathbf{z}^{(2)}_{\text{pos}}$ and similarity between $\mathbf{z}$ and $\mathbf{z}_{\text{neg}}$. Our contrastive loss for transformation latent factor can be written as:
$L_{\text{con(z)}}=L_{\text{dist}}(\mathbf{z}^{(1)}_{\text{pos}},\mathbf{z}^{(2)}_{\text{pos}})+L_{\text{sim}}(\mathbf{z},\mathbf{z}_{\text{neg}})$ (4)
Similar to $L_{\text{con}(\mathbf{c})}$, we use mean absolute cosine distance and mean absolute cosine similarity to implement $L_{\text{dist}}$ and $L_{\text{sim}}$, respectively.
Objective Function: In summary, we train the encoder and decoder networks by simultaneously minimizing the loss components. Our overall objective function to minimize is:
$L=L_{\text{vae}}+L_{\text{con(c)}}+L_{\text{con(z)}}$ (5)
4 Experiments & Results
Table 1: Quantitative Results of unsupervised Content-Transformation Disentangling Methods on MNIST and LineMod datasets. This evaluation only shows the disentanglement performance w.r.t ground truth content factors. We report mean $\pm$ std. dev. over model training by setting 3 different random seeds. Here, for Harmony and SpatialVAE, the transformation latent factor is restricted to dimension 3. For the other methods, there is no such restriction on the transformation latent factor. For SimCLR, (Disc) means discriminative model (without any reconstruction), and (Gen) means Generative model (with reconstruction).
MNIST LineMod
Method $D_{\text{score}}(c|c)(\uparrow)$ $D_{\text{score}}(c|s)(\downarrow)$ $D_{\text{score}}(c|c)(\uparrow)$ $D_{\text{score}}(c|s)(\downarrow)$
SpatialVAE [2] 0.81 $\pm$ 0.01 0.28 $\pm$ 0.02 0.95 $\pm$ 0.02 0.32 $\pm$ 0.01
Harmony [4] 0.82 $\pm$ 0.01 0.31 $\pm$ 0.02 0.90 $\pm$ 0.01 0.56 $\pm$ 0.02
SimCLR (Disc) with rotation [15] 0.58 $\pm$ 0.01 0.60 $\pm$ 0.01 0.62 $\pm$ 0.01 0.40 $\pm$ 0.01
SimCLR (Gen) with rotation [15] 0.53 $\pm$ 0.01 0.67 $\pm$ 0.02 0.61 $\pm$ 0.02 0.79 $\pm$ 0.01
VAE with two latent space 0.63 $\pm$ 0.00 0.63 $\pm$ 0.00 0.87 $\pm$ 0.00 0.87 $\pm$ 0.00
VITAE [1] 0.77 $\pm$ 0.00 0.32 $\pm$ 0.00 0.92 $\pm$ 0.00 0.90 $\pm$ 0.00
DualContrast (w/o $L_{\text{con(c)}}$) 0.87 $\pm$ 0.02 0.21 $\pm$ 0.02 0.86 $\pm$ 0.01 0.31 $\pm$ 0.02
DualContrast (w/o $L_{\text{con(z)}}$) 0.79 $\pm$ 0.04 0.85 $\pm$ 0.02 0.79 $\pm$ 0.03 0.86 $\pm$ 0.03
DualContrast 0.89 $\pm$ 0.01 0.31 $\pm$ 0.02 0.95 $\pm$ 0.01 0.48 $\pm$ 0.02
Benchmark Datasets: We report results on MNIST [34], LineMod [35], Starmen [36], and a realistically simulated protein subtomogram dataset. Only LineMod features RGB images among these datasets, while the other three consist of grayscale images. The subtomogram is a 3D volumetric dataset, whereas the others are 2D images.
Baselines: We used the unsupervised content-transformation disentanglement methods, SpatialVAE [2], Harmony [4], VITAE [1] with explicit parameterization and InfoNCE loss based standard self-supervised learning method with rotation augmentation [15] as baseline approaches. Several generic disentangled representation learning methods, such as beta-VAE, Factor-VAE, beta-TCVAE, DIP-VAE, etc., are available that do not aim to disentangle specific factors such as content or transformation in the data but rather aim to disentangle all factors of variation. However, previous studies [2, 1] have shown that such methods perform worse than methods specifically aiming to disentangle content and transformation. Therefore, we did not use these methods as baselines in our experiments.
Implementation Details: We implemented the encoder using a Convolutional Neural Network (CNN) and the decoder using a Fully Connected Network (FCN). Our experiments used the same latent dimension for content and transformation factors, except for Harmony [4] and SpatialVAE [2], where a specific dimension needs to be used for transformation latent factors. We train our models for $200$ epochs with a learning rate of $0.0001$ in NVIDIA RTX A500 GPUs and AMD Raedon GPUs. We optimize the model parameters with Adam optimizer.
Evalutation Metrics: We assess DualContrast based on two criteria: (1) the quantitative disentanglement of content and transformation factors and (2) the qualitative capability of image generation through content-transformation transfer.
Figure 2: Qualitative content-transformation transfer based image generation results for MNIST. (a), (b), (c), and (d) shows the results with Harmony [4], Spatial-VAE [2], VITAE [1], and DualContrast respectively. Erroneous generations (both in terms of content and transformation) are marked within red boxes.
For (1), there exists several metrics, e.g., MIG score, $D_{\text{score}}$, etc. Locatello et al [37] demonstrated that these metrics are highly correlated. To this end, similar to [1], we only use $D_{\text{score}}$ to measure the disentanglement. $D_{\text{score}}$ is simply a measurement of how well the ground truth factors can be predicted from the corresponding latent factors. In our scenario, there are four such quantities - (a) predictivity of content labels given content factor $D_{\text{score}}(c|c)$ or transformation factor $D_{\text{score}}(c|s)$ (b) predictivity of transformation labels given content factor $D_{\text{score}}(s|c)$ or transformation factor $D_{\text{score}}(s|s)$. For MNIST and LineMod, we do not have any transformation labels, so we only reported values for (a). Similar to [15], we use linear Logistic Regression Classifiers to measure predictivity ($D_{\text{score}}$) across all our experiments.
Figure 3: tSNE embedding plots of content latent factor learned by the unsupervised content-style disentanglement methods. (a), (b), (c), and (d) shows the results for Harmony [4], Spatial-VAE [2], C-VITAE [1], and DualContrast respectively. Overall, DualContrast shows superior performance.
The quantitative measurement, particularly when transformation labels are not available, might be misleading. To this end, we perform qualitative evaluation by generating images while transferring content and transformation from original images from the dataset. We take a set of $K$ images $\mathbf{x}$ and create a $K\times K$ grid of $K^{2}$ images using the different combinations of content and transformation factors from those $K$ images. In the grid, we keep the content factor and transformation factor consistent across the columns and the rows respectively. In other words, the image at position $(i,j)$ in the grid is generated using the content of image $x_{j}$ and transformation of image $x_{i}$. For the protein subtomogram dataset, we also provide qualitative results through embedding visualization (Figure 8).
4.1 DualContrast disentangles writing style from MNIST handwritten digit images
Similar to [4, 2, 1], we start our experiments with the MNIST dataset of handwritten digits. We performed a quantitative comparison between DualContrast and the baseline approaches, as shown in Table 1. We observed that DualContrast achieves higher $D_{score}(c|c)$ compared to other explicitly parameterized approaches [4, 2, 1].
Figure 4: Qualitative Results of Unsupervised Content-Transformation Disentanglement on LineMod obtained by VITAE[1] (a) and DualContrast (b), respectively.
However, since transformation labels are not present for MNIST, relying entirely on the content prediction performance for disentanglement might be misleading. For example, the degenerate solution achieves even lower $D_{score}(c|s)$ as transformation does not represent anything. To this end, we report qualitative results for baseline methods and our approach DualContrast (Fig. 2). In generating images with varying content and styles, we observe that Harmony, SpatialVAE, and C-VITAE generate many images with erroneous content and styles. On the other hand, DualContrast produces a few minor errors and no extreme mistakes. We also include tSNE embedding of the content factors inferred by the models on the MNIST test dataset associated with class labels (Figure 3). On the embedding space, DualContrast clearly shows superior clustering performance. All of these suggest better content-transformation disentanglement with our approach than existing explicit parameterization approaches.
4.2 DualContrast Disentangles ViewPoint from Linemod Object Dataset
Figure 5: Qualitative Results of Unsupervised Content-Transformation Disentanglement on Starmen (a) and (b) shows the results obtained with VITAE [1] and DualContrast, respectively.
LineMod is an RGB object recognition and 3D pose estimation dataset [35] with 15 unique object types. It contains 15 unique objects: ‘ape’, ‘bench vise’, ‘bowl’, ‘cam’, ‘can’, ‘cat’, ‘cup’, ‘driller’, ‘duck’, ‘eggbox’, ‘glue’, ‘hole puncher’, ‘iron’, ‘lamp’ and ‘phone’, photographed in a highly cluttered environment. We use a synthetic version of the dataset [38], which has the same objects rendered under different viewpoints. The dataset is publicly available at this url. The dataset has $,31$ 313 images per object category. We used $1,000$ images per category for training and the remaining for testing.
We evaluated the disentanglement scores $D_{\text{score}}$ (Table 1) for DualContrast and baseline approaches, where DualContrast clearly shows the best performance. We used the object identities as content labels during the evaluation.
We also performed qualitative measurement with image transfer and depicted the result of DualContrast and its most relevant (in terms of the generality of transformations) VITAE in Fig 4. We observe that VITAE distorts the images significantly while performing image generation with content-style transfer. On the other hand, DualContrast does not face such issues. Moreover, DualContrast can disentangle different viewpoints of the objects as the transformation factor using only 2D rotation to create contrastive pairs in its model.
4.3 DualContrast disentangles Human Deformation from starmen shapes dataset
Starmen is a 2D synthetic shape dataset of the deformable human body [36] consisting of $10,000$ shape-focused images, which is publicly available at https://zenodo.org/records/5081988 under CC-BY license. Each starmen image is a grayscale image of size $64\times 64$. They are generated by deforming five control points (two on arms, two on legs, one on head) on template human shapes.
We randomly took $6,000$ starmen images to train our models and kept the remaining $4,000$ for testing. We report the qualitative results by VITAE [1] and DualContrast in Figure 5. From the image generation results with the content-transformation transfer, we can observe that DualContrast disentangles the deformation of starmen shapes as transformation and the identity (based on leg thickness) as content, providing a smooth image generation result. On the other hand, VITAE [1] does not perform such meaningful content-transformation disentanglement, generating implausible results.
4.4 DualContrast Disentangles Protein Identity from states in Cryo-ET Subtomograms
Figure 6: (a) Isosurface visualization of Nucleosome Density Maps. Each density map slightly varies in terms of conformation. They are associated with the PDB IDs in the figure. (b) 3D Slice by slice visualization of several subtomograms of the corresponding nucleosome density maps.
We created a realistic simulated cryo-ET subtomogram dataset of 18,000 subtomograms of size $32^{3}$. The dataset consists of 3 protein classes of similar sizes- Nucleosomes, Sars-Cov-2 spike protein, and Fatty Acid Synthase Unit. These proteins are significantly different in terms of their composition, which determines their different identities. Moreover, structures for all of these three types of proteins have been resolved in cellular cryo-ET [39, 30, 40], which makes it feasible to use them for our study. Moreover, cellular cryo-ET is the primary method to capture these proteins inside the cells in their native state [41].
For each protein class, we collected 6 different protein structures from the RCSB PDB website [42]. RCSB PDB is a web server containing the structure of millions of proteins. For nucleosomes, we collected PDB IDs ‘2pyo’, ‘7kbe’, ‘7pex’, ‘7pey’, ‘7xzy’, and ‘7y00’. All of these are different conformations of nucleosomes that slightly vary in composition (Figure LABEL:fig:supp:nuc-dmaps). For sars-cov-2 spike proteins, we collected PDB IDs ‘6vxx’, ‘6vyb’, ‘6xr8’, ‘6xra’, ‘6zox’, and ‘6zp0’. Among them, only ‘6xra’ shows much variation in structure from the other ones, and the rest of the PDB IDs are very similar in structure. For Fatty Acid Synthase (FAS) Unit, we used PDB IDs ‘8prv’, ‘8ps1’, ‘8ps9’, ‘8psj’, ‘8psm’, ‘8psp’. They also vary very slightly in terms of the structure.
After collecting these 18 PDB structures as PDB files, we used EMAN PDB2MRC [43] to create density maps (as MRC file extension) from the PDB files. We create density maps of size $32^{3}$ with $1$ nm resolution. We then randomly rotate and translate each density map and create $1000$ such copies. We then convolve the density maps with CTF with CTF parameters common in experimental datasets (Defocus -5 nm, Spherical Abberation 1.7, Voltage 300 kV). Afterward, we add noise to the convolved density maps so that the SNR is close to $0.1$. Thus, we prepare $18,000$ realistic subtomograms with 3 different protein identities, each with 6 different conformations. We uploaded the entire dataset anonymously at https://zenodo.org/records/11244440 under CC-BY-SA license. 3D slice-by-slice visualizations for sample subtomograms for nucleosomes, spike proteins, and FAS units are provided in Figure 6.
We trained DualContrast and baseline approaches against the protein subtomogram dataset. We could not train VITAE [1] on subtomogram datasets since it did not define any transformation for 3D data. Designing CPAB transformation for 3D data by ourselves was challenging. Among the other baseline approaches, spatialVAE could not distinguish the protein identities with high heterogeneity at all, which is evident by its embedding space UMAP (Figure 7). Only Harmony provided somewhat plausible results apart from DualContrast.
Figure 7: UMAP of content factor learned by SpatialVAE [2] on our subtomogram dataset. The protein identities are indistinguishable from the embedding space.
We observed that the DualContrast content factor perfectly clustered the three protein identities from the dataset (GMM clustering accuracy $0.99$), profoundly disentangling the transformations (Figure 8(b)). We do not see such perfect clustering for Harmony (Figure 8(a)); rather, protein identities with very different compositions get mixed up due to their conformational variation (GMM clustering accuracy $0.70$). We further observe that DualContrast transformation embedding clusters have similar nucleosome conformations, showing disentanglement of conformation from protein identity. On the other hand, in Harmony, the transformation factor only represents rotation and translation by design and can not capture nucleosome conformations. Moreover, even its content embedding for manually selected (not through automatic clustering) nucleosome subtomograms can not show good clustering of conformations (Figure 8(b)).
Figure 8: Disentanglement of composition and conformations in cellular subtomogram dataset. (a) UMAP embedding of the content factor of all subtomograms (top) and only nucleosome subtomograms (bottom) with Harmony [4] (b) UMAP embedding of the content factor (top) and transformation factor (bottom) obtained by DualContrast.
Ablation Study: To evaluate the individual contribution of the contrastive losses, we conduct both quantitative and qualitative ablation analyses of DualContrast. We trained (1) DualContrast without any contrastive loss, which is basically a VAE with two latent spaces, (2) DualContrast with only $L=L_{\text{VAE}}+L_{\text{con (z)}}$, and (3) $L=L_{\text{VAE}}+L_{\text{con (c)}}$. We qualitatively and quantitatively evaluated each model. We show the quantitative results in Table 1 and qualitative results of MNIST in Figure 9.
Figure 9: Content-transformation transfer results from ablation analysis. (a) and (b) shows the results when the model is trained with $L_{\text{VAE}}+L_{\text{con (z)}}$ and $L_{\text{VAE}}+L_{\text{con (c)}}$ respectively.
For (1), the $D_{\text{score}}$ is almost the same for both content and transformation factors, indicating equal predictivity of the digit classes by both factors. This is obvious given that the model has no inductive bias to make different factors capture different information. In model (2), using contrastive loss w.r.t. only transformation factor $z$ makes it uninformative of the data. Thus, it provides a small $D(c|z)$ score as desired, but the changing $z$ does not affect the image generation (Figure 9). On the other hand, in the model (3), using contrastive loss w.r.t. only content factor $c$ gives a less informative content factor, a lower $D(c|c)$ score, and a higher $D(c|z)$ score, contrary to what is desired. These results indicate contrastive loss w.r.t to both factors is crucial to get the desired disentanglement.
5 Limitation
Similar to existing content-transformation disentanglement methods [2, 4, 1, 8, 12], DualContrast is applicable to data samples with single objects. This is reasonable, given content and transformation can not be defined for data with multiple objects. Moreover, we limited our study only to shape-focused image datasets, motivated by the problem of conformation disentanglement of protein subtomograms. Consequently, we only used rotation to create contrastive pairs, which worked pretty well in our cases. For image datasets where appearance or illumination plays a crucial role, using only rotation to create contrastive pairs may not work well. However, disentanglement in natural image datasets is a different scenario that is outside of the scope of this paper.
6 Conclusion
This work focuses on a challenging setting of the unsupervised content-transformation disentanglement problem in shape-focused image datasets, where the transformation latent factor is not explicitly parameterized. To tackle this problem, we propose a novel method termed DualContrast. DualContrast employs generative modeling with a novel contrastive learning strategy that creates positive and negative pairs for content and transformation latent factors. Our extensive experiments show DualContrast’s effective disentanglement of challenging transformation factors across various shape-based image datasets; including simulated cellular subtomograms, where it solved the unexplored problem of isolation of protein conformations from protein identities, as proof of principle.
7 Acknowledgement
This work was supported in part by U.S. NIH grants R01GM134020 and P41GM103712, NSF grants DBI-1949629, DBI-2238093, IIS-2007595, IIS-2211597, and MCB-2205148. This work was supported in part by Oracle Cloud credits and related resources provided by Oracle for Research, and the computational resources support from AMD HPC Fund. MRU were supported in part by a fellowship from CMU CMLH.

### v2

@@
 Abstract
-Unsupervised disentanglement of content and transformation has recently drawn much research, given their efficacy in solving downstream unsupervised tasks like clustering, alignment, and shape analysis.
-This problem is particularly important for analyzing shape-focused real-world scientific image datasets, given their significant relevance to downstream tasks.
-The existing works address the problem by explicitly parameterizing the transformation factors, significantly reducing their expressiveness.
+Unsupervised disentanglement of content and transformation is significantly important for analyzing shape focused scientific image datasets, given their efficacy in solving downstream image-based shape-analyses tasks.
+The existing relevant works address the problem by explicitly parameterizing the transformation latent codes in a generative model, significantly reducing their expressiveness.
 Moreover, they are not applicable in cases where transformations can not be readily parametrized.
-An alternative to such explicit approaches is self-supervised methods with data augmentation, which implicitly disentangles transformations and content.
-We demonstrate that the existing self-supervised methods with data augmentation result in the poor disentanglement of content and transformations in real-world scenarios.
-Therefore, we developed a novel self-supervised method, DualContrast, specifically for unsupervised disentanglement of content and transformations in shape-focused image datasets.
-Our extensive experiments showcase the superiority of DualContrast over existing self-supervised and explicit parameterization approaches.
-We leveraged DualContrast to disentangle protein identities and protein conformations in cellular 3D protein images.
-Moreover, we also disentangled transformations in MNIST, viewpoint in the Linemod Object dataset, and human movement deformation in the Starmen dataset as transformations using DualContrast.
-Conceptualization of content-transformation disentanglement (a) and comparison between existing explicit (b) and our implicit (c) unsupervised methods.
-Both of these methods optimize content and transformation latent spaces while reconstructing $\mathbf{x}^{\prime}$ from $\mathbf{x}$.
-However, explicit methods use the transformation space to infer a fixed parameter set $\gamma$ and use it to transform the feature map/image $\mathbf{x}_{c}$ generated by conditioning the decoder on content space.
-On the contrary, our implicit method conditions the decoder on both content and transformation to generate image $\mathbf{x}^{\prime}$ and does not restrict the transformation space to a fixed set of parameters.
+An alternative to such explicit approaches is contrastive methods with data augmentation, which implicitly disentangles transformations and content.
+However, the existing contrastive strategies are insufficient to this end.
+Therefore, we developed a novel contrastive method with generative modeling, DualContrast, specifically for unsupervised disentanglement of content and transformations in shape focused image datasets.
+DualContrast creates positive and negative pairs for content and transformation from data and latent spaces.
+Our extensive experiments showcase the efficacy of DualContrast over existing self-supervised and explicit parameterization approaches.
+With DualContrast, we disentangled protein composition and conformations in cellular 3D protein images, which was unattainable with existing disentanglement approaches.
 1 Introduction
-Figure 1: Our proposed contrastive learning-based unsupervised content-transformation disentanglement pipeline.
-(a) The variational inference of content and transformation factors with $L_{\text{VAE}}$.
-(b,c) The contrastive strategy to create positive and negative pairs w.r.t.
-content (b) and transformation factors (c), respectively.
-DualContrast combines the three parts and is trained end-to-end.
+Figure 1: (a) The concept of content-transformation disentanglement, whereas changing the content changes the protein identity, and changing transformation changes the state of the protein.
+Explicit methods (b) use the transformation space to infer a fixed parameter set, whereas our implicit method (c) do not restrict the transformations to a fixed set of parameters.
+Figure uses toy protein images for visualization, they are not used for experiments in this form.
 In many real-world image datasets, particularly in scientific imaging domains, the object shapes being visualized may undergo multiple transformations [1, 2].
-These shape-focused image datasets can thus be regarded as samples generated from two independent factors, one representing the semantic attribute termed content and the other representing the transformations.
+These shape focused image datasets can thus be regarded as samples generated from two independent factors, one representing the semantic attribute termed content and the other representing the transformations.
 Content is regarded as the form of an object being visualized in the image [1] that is unaltered after applying any nuisance transformation.
 On the other hand, transformation dictates the specific state or realization of that form of the object.
-Taking proteins inside the cells as an example, changing the transformation factor changes the conformation of the protein where the protein identity (i.e., protein composition) does not change (Figure DualContrast: Unsupervised Disentangling of Content and Transformations with Implicit Parameterization).
+Taking proteins inside the cells as an example, changing the transformation factor changes the conformation of the protein where the protein identity (i.e.
+, protein composition) does not change (Fig.
+1).
 On the other hand, changing the content factor is analogous to changing the proteins from one identity to another through compositional change.
-Such a phenomenon holds for many real-world shape-focused datasets.
-Disentangling content and transformation factors of images can significantly facilitate several downstream tasks, including clustering [3], alignment [4, 2], zero-shot transfer [5], bias removal [6, 7], cross-domain retrieval [8, 9], shape analysis [10, 11, 4, 2, 3], etc.
+Such a phenomenon holds for many shape focused image datasets.
+Disentangling content and transformation factors of images can significantly facilitate several downstream shape analysis tasks, including shape clustering [3], alignment [4, 2], zero-shot transfer [5], bias removal [6, 7], cross-domain retrieval [8, 9], etc.
 [10].
@@
 Only a few works [1, 2, 4] have addressed the problem in a completely unsupervised setting.
-Among them, VITAE [1] performed explicit parameterization of transformation factors as diffeomorphic transformations and achieved noteworthy disentanglement of content and transformation in several image datasets.
-SpatialVAE [2] and Harmony [4] focus on representing known transformation types as transformation factors.
-SpatialVAE [2] performed disentanglement of two-dimensional translation and rotation, whereas Harmony [4] demonstrated disentanglement of a broader set of transformations that can be explicitly parameterized.
-Though these unsupervised explicit parameterization methods have demonstrated several successes, they have several major limitations in general: (1) These method uses the transformation factor to generate only a few parameterized transformations.
-In reality, many transformations do not have a well-defined parameterized form,   e.g., protein conformation changes, viewpoint change (LineMod), human deformation during movement (Starmen), etc.
+Among them, [1] performed explicit parameterization of transformation codes as diffeomorphic transformations and achieved noteworthy disentanglement of content and transformation in several image datasets.
+[2] and [4] focus on representing known transformation types as transformation codes.
+[2] performed disentanglement of two-dimensional translation and rotation, whereas [4] demonstrated disentanglement of a broader set of transformations that can be explicitly parameterized.
+Though these unsupervised explicit parameterization methods have demonstrated several successes, they have several major limitations in general: (1) These method uses the transformation codes to generate only a few parameterized transformations.
+In reality, many transformations do not have a well-defined parameterized form,   e.g., protein conformation changes, viewpoint change (LineMod), etc.
 These explicit methods can not disentangle such transformation by design.
-(2) They use Spatial Transformer Networks (STN) [13] for inferring the transformation factor, which requires a continuous parameterization of the transformation to be disentangled.
-The continuity of many transformations (e.g., SO(3) rotation) in neural networks is often a concern [14].
-Moreover, (3) these methods often infer spurious contents during disentanglement in real-world datasets(Fig.
+(2) They use Spatial Transformer Networks (STN) [13] for inferring the transformation codes, which requires a continuous parameterization of the transformation to be disentangled.
+The continuity of many transformations (e.g.
+, SO(3) rotation) in neural networks is often a concern [14].
+Moreover, (3) these methods infer spurious contents during disentanglement in real scenarios where transformations not predefined are present in the dataset (Fig.
+3).
+Due to these limitations, an unsupervised content-transformation disentangling method without explicit transformation representation is much desired.
+Consequently, in this work, we focused on the more generalized and unique setting of unsupervised content-transformation disentanglement without any explicit parameterization of the transformation code.
+Previously, a theoretical study by [15] showed that standard Contrastive Learning (CL) methods, e.g., SimCLR [16], can disentangle content from style, where style can be regarded as the transformations used for data augmentation.
+However, these popular CL methods [16] have rarely achieved any disentanglement in practice [6], and their ability in disentangling transformations other than those used for augmentation is not well-explored.
+Moreover, we found that SimCLR with geometric data augmentation can not disentangle any transformation well in our shape focused image datasets of interest (Table 1).
+Figure 2: Our proposed contrastive learning-based unsupervised content-transformation disentanglement pipeline.
+(a) The variational inference of content and transformation codes with $L_{\text{VAE}}$.
+(b) Contrastive pair creation strategy for content and transformation codes.
+The process is delineated in the bottom panel.
+Additional visualization available in Appendix Fig.
+9.
+(c) Contrastive losses.
+In DualContrast, the contrastive pair creations and reconstruction happen simultaneously, and the encoder and decoder network are optimized with both contrastive and reconstruction losses in each iteration.
+To this end, we develop a novel generative contrastive learning method for implicit content-transformation disentanglement, termed “DualContrast” (Fig.
 2).
-Due to these limitations, an unsupervised content-transformation disentangling method without explicit transformation representation is much desired.
-Consequently, in this work, we focused on the more generalized setting of unsupervised content-transformation disentanglement without any explicit parameterization of the transformation factor.
-Given the theoretical study by Von et al.
-[15] showing the implicit disentanglement ability of standard Contrastive Learning (CL) methods, e.g., SimCLR [16], to disentangle content from style, a plausible solution could be using SimCLR with geometric data augmentation for our implicit content-transformation disentanglement task.
-However, these popular CL methods [16] have rarely achieved disentanglement in practice [6], and we also found that SimCLR with geometric data augmentation performs poorly in our shape-focused image datasets of interest (Table 1).
-To this end, we develop a novel contrastive learning method for implicit content-transformation disentanglement, termed “DualContrast”.
-Similar to previous content-transformation disentangling methods [17, 2, 1, 4], it is a deep generative model based on a variational autoencoder (VAE) architecture.
-However, unlike previous methods, it does not require labels or explicit parameterization of any latent factor.
-Instead, it uses a novel contrastive strategy that creates “positive” and “negative” contrastive pairs w.r.t.
-both content and transformation latent factor.
-We create negative pairs w.r.t.
-transformation factor and positive pairs w.r.t.
-content factor by simply rotating the samples.
+Unlike existing contrastive strategies, it creates “positive” and “negative” contrastive pairs w.r.t.
+both content and transformation latent codes.
+We found that several shape transformations in image datasets lie closely in the normally distributed latent space that is designed to capture in-plane rotation information in a contrastive manner.
+Consequently, we create negative pairs w.r.t.
+transformation and positive pairs w.r.t.
+content by simply rotating the samples.
+Without prior knowledge, creating positive pairs w.r.t.
+the transformation codes directly from the input samples is not possible.
+To this end, we generate samples from an identical distribution of the transformation representations in the embedding space and treat them as positive pairs for the transformation codes.
+Given our observation, this encourages the disentanglement of several shape transformations other than rotation in our dataset.
 In addition, we randomly permute the samples in the batch to create a negative pair for a batch of samples w.r.t.
-content latent factor.
-Without prior knowledge, creating positive pairs w.r.t.
-the transformation latent factor from the input samples is not possible.
-To this end, we generate samples from an identical distribution of the transformation representations in the embedding space and treat them as positive pairs for the transformation latent factor.
+content.
 Upon creating both “positive” and “negative” pairs w.r.t.
-both factors, we penalize similarity between negative pairs and distance between positive pairs for each factor.
-In DualContrast, we combine these contrastive losses with VAE reconstruction loss and train the model in an end-to-end manner.
-With such a strategy, by just performing simple rotation to create contrastive pairs, DualContrast disentangles a much wider range of more complicated transformations, e.g., human deformation in starmen, viewpoint in Linemod, the writing style in MNIST, etc.
-With DualContrast, we, for the first time, solved the scientific problem of disentangling protein identities from protein conformations in simulated 3D protein images (Figure 6) inside the cell, known as cellular cryo-ET subtomograms.
-DualContrast could successfully disentangle the protein identities as content and protein conformations as transformation factors in simulated subtomograms, which is not achievable with previous methods.
+both codes, we penalize similarity between negative pairs and distance between positive pairs for each code.
+Finally, we incorporate these contrastive losses in a Variational Autoencoder (VAE) based architecture with VAE reconstruction loss and train the model in an end-to-end manner.
+With such a strategy, by just performing simple rotation to create contrastive pairs, DualContrast disentangles a much wider range of transformations caussing subtle changes in the pixel space,   e.g., viewpoint in Linemod, several writing styles in MNIST, etc.
+With DualContrast, we, for the first time, solved the scientific problem of disentangling protein identities from protein conformations (defined as subtle changes) in simulated 3D protein images (Fig.
+5 and Fig.
+6) inside the cell, known as cellular cryo-ET subtomograms.
+DualContrast could successfully disentangle the protein identities as content and protein conformations as transformation in simulated subtomograms, which is not achievable with previous methods.
 Overall, our qualitative and quantitative experimental results show the efficacy of DualContrast in isolating content from transformation compared to existing relevant methods.
 A summary of our contributions is as follows:
-  • We work on a challenging setting of learning unsupervised disentangled representations of content and transformation in shape-focused images by removing the requirement of explicit parameterization.
-  • We develop a novel contrastive generative modeling strategy that creates positive-negative pairs for both content and transformation latent factors.
-We introduced a novel way of designing contrastive pairs for transformation factors.
-  • We show that DualContrast is highly effective in disentangling a wide range of transformations from the content in various shape-focused image datasets by only using simple rotation for creating contrastive pairs.
-  • Leveraging DualContrast, we, for the first time, disentangled protein identities (i.e.composition) from protein conformations in simulated 3D cellular cryo-ET subtomograms, as a proof of principle.
+  • We develop a novel contrastive generative modeling strategy that creates positive-negative pairs for both content and transformation latent codes.
+We introduced a novel way of designing contrastive pairs for transformation from both data and latent space.
+  • We show that DualContrast is effective in disentangling a wide range of transformations that causes subtle changes in pixel space, from the content in various shape focused image datasets.
+  • Leveraging DualContrast, we, for the first time, disentangled protein identities (i.e.
+composition) from protein conformations in simulated 3D cellular cryo-ET subtomograms, as a proof of principle.
+DualContrast is the first method that can identify distinct conformations of proteins from protein mixture subtomogram datsets.
 2 Related Works
-Disentangled Representation Learning: Our method can be regarded as a disentangled representation learning method that disentangles generative factors of variation as content-specific and transformation-specific factors.
-There exist many generic disentangled representation learning approaches [18, 17, 19, 20, 21, 22] where the factors of variations are interpreted by traversing through the learned factors and reconstructing samples.
-However, these methods do not perform very well in specifically disentangling content and transformation factors [1].
-To this end, specialized methods that deal with such two factors have been introduced recently.
-Many of them [23, 12, 8] are partially or weakly supervised and use labels associated with one of the latent factors to train their models.
-However, these methods are not applicable when such supervision is unavailable beforehand, which is often the case for many real-life datasets.
-There are a few [1, 4, 2, 15] approaches that deal with the unsupervised disentanglement of content and transformation.
-However, these methods perform explicit parameterization of the transformation factor as affine or some specific parameterized transformation.
-In contrast, our method does not impose any explicit parameterization on the transformation latent factor and does not face the issues of the explicit methods (Section 1).
+Unsupervised Content-Transformation Disentanglement: There are a few [1, 4, 2, 15] approaches that deal with the unsupervised disentanglement of content and transformation.
+However, these methods perform explicit parameterization of the transformation codes as affine or some specific parameterized transformation.
+In contrast, our method does not impose any explicit parameterization on the transformation latent codes and does not face the issues of the explicit methods (Section 1).
 Contrastive Learning based Disentanglement: Contrastive learning is the primary building block of our method.
-Contrastive Learning with data augmentation with existing approaches, e.g., InfoNCE loss-based SimCLR [16], etc., has been previously used by [15, 8] to isolate content from style.
-Though in several scenarios, style in shape focused images can be referred to as transformations, these works do not specifically focus on shape-focused real-world images.
+Contrastive Learning with data augmentation with existing approaches, e.g.
+, InfoNCE loss-based SimCLR [16], etc., has been previously used by [15, 8] to isolate content from style.
+Though in several scenarios, style in shape focused images can be referred to as transformations, these works do not specifically focus on shape focused real-world images.
 Moreover, our experiments demonstrate that SimCLR with rotation augmentation in a standard encoder-decoder framework leads to very poor disentanglement of content and transformation (Table 1).
-Moreover, [15] only creates positive pairs w.r.t content and negative pair w.r.t.
+Also, [15] only creates positive pairs w.r.t content and negative pair w.r.t.
 style with data augmentation in SimCLR.
 We have observed this scheme does not work well in our scenario.
-Unlike these works, our approach uses a novel strategy to create “negative" and “positive" pairs for both content and transformation of latent factors without any InfoNCE loss.
+Unlike these works, our approach uses a novel strategy to create “negative” and “positive” pairs for both content and transformation of latent codes without any InfoNCE loss.
 Unsupervised Content-style disentanglement: Apart from [15], several methods [6, 17, 18, 19, 20, 21] exist that perform unsupervised content-style disentanglement, focusing on natural images.
-These methods focus on the controllable generation of realistic natural images, mostly using ImageNet pretrained models.
-Unlike these methods, our work primarily focuses on disentangling content and transformation in shape-focused image datasets.
+Unlike these methods, our work primarily focuses on disentangling content and transformation in shape focused image datasets.
 Moreover, we do not depend on any ImageNet pretrained models as our images of interest differ greatly from the natural images of ImageNet.
-Protein Composition-Conformation Disentanglement: One of the major contributions of our work is disentangling protein identity characterized by large compositional variability as the content factor and protein states characterized by conformational variability or subtle compositional variability as transformation factor from 3D cellular cryo-electron tomography (cryo-ET) subimages or subtomograms.
+The very recent work by [6] assumes access to the style factors to disentangle that style from content in feature outputs from pretrained models.
+Unlike this work, we do not assume access to the transformations beforehand for disentanglement.
+Protein Composition-Conformation Disentanglement: One of the major contributions of our work is disentangling protein identity characterized by large compositional variability as the content codes and protein states characterized by conformational variability or subtle compositional variability as transformation codes from 3D cellular cryo-electron tomography (cryo-ET) subimages or subtomograms.
 Previously, Harmony [4] was used to disentangle large compositional variability in cellular subtomograms.
@@
 SpatialVAE [2] disentangled subtle conformational variability from 2D cryo-EM images.
-There exists several works on single particle cryo-EM and cryo-ET reconstruction, e.g., cryoDRGN2 [31], cryoFIRE [3], cryoAI [32], etc.
-that performs amortized inference of transformation ($SO(3)\times d^{2}$) and latent space representing content.
-However, these works mainly focus on 2D-to-3D reconstruction instead of content-transformation disentanglement.
 Ours is the first work to deal with both large and subtle compositional variability in two different latent spaces for 3D cryo-ET subtomograms.
+Further discussions on the related works can be found in Appendix. Section A.1.
 3 Method
@@
 In this paper, we distinguish between content and transformation with respect to a family of transformations $\mathcal{T}:\mathcal{X}\rightarrow\mathcal{X}$ such that the following two conditions hold:
-  • Condition 1: $\forall$ $T\in\mathcal{T}$, s.t., for any $\mathbf{x}\in\mathcal{X}$, $h_{c}(T(\mathbf{x}))=h_{c}(\mathbf{x})$.
-  • Condition 2: $\exists$ $T\in\mathcal{T}$, s.t., for any $\mathbf{x}^{(1)},\mathbf{x}^{(2)}\in\mathcal{X}$, $h_{z}(T(\mathbf{x}^{(1)}))=h_{z}(\mathbf{x}^{(2)})$.
+  • Condition 1: $\forall$ $T\in\mathcal{T}$ and $\forall$ $\mathbf{x}\in\mathcal{X}$, $h_{c}(T(\mathbf{x}))=h_{c}(\mathbf{x})$.
+  • Condition 2: $\exists$ $T\in\mathcal{T}$ and $\forall$ $\mathbf{x}^{(1)},\mathbf{x}^{(2)}\in\mathcal{X}$, $h_{z}(T(\mathbf{x}^{(1)}))=h_{z}(\mathbf{x}^{(2)})$.
 Condition 1 defines content space which is invariant w.r.t.
@@
 Thus it defines transformation space which is informative of $\mathcal{T}$.
-For content-transformation disentangling, changes in $\mathcal{C}$ must be unaffected by changes in $\mathcal{Z}$ and vice versa.
+For content-transformation (c-z) disentangling, changes in $\mathcal{C}$ must be unaffected by changes in $\mathcal{Z}$ and vice versa.
 In our problem setting, no labels associated with content space $C$ or transformation space $Z$ are known.
@@
 3.2 Method Overview and Notation
-Our method is built upon a variational auto-encoder (VAE) based architecture (Figure 1), where we perform inference of the content and transformation factors in the data.
+Our method is built upon a variational auto-encoder (VAE) based architecture (Fig.
+2), where we perform inference of the content and transformation factors in the data.
 Consider, a sample space $\mathcal{X}$, two latent distribution spaces $\Phi_{Z}$ and $\Phi_{C}$, two latent spaces $\mathcal{Z}$ and $\mathcal{C}$.
@@
 These parameters are then decoded using a decoder network $p_{\theta}:\mathcal{C}\times\mathcal{Z}\rightarrow\mathcal{X}$ to produce $x_{\text{recon}}=p_{\theta}(\mathbf{c},\mathbf{z})$.
-3.3 Variational Inference of Content and Transformation Factors
+3.3 Variational Inference of Content and Transformation Codes
 Similar to a variational auto-encoder (VAE) setting, the encoder and decoder network parameters are simultaneously optimized to maximize the variational lower bound to the likelihood $p(\mathbf{x})$ called the evidence lower bound (ELBO).
@@
 In short, we write the KL terms as $L_{\text{KL}(c)}$ and $L_{\text{KL}(z)}$.
-To regularize the effect of prior on the final objective separately for two latent factors, we scale the two losses with two hyperparameters $\gamma_{c}$ and $\gamma_{z}$.
+To regularize the effect of prior on the final objective separately for two latent codes, we scale the two losses with two hyperparameters $\gamma_{c}$ and $\gamma_{z}$.
 Our loss function for variational inference becomes:
 $L_{\text{vae}}=L_{\text{rec}}+\gamma_{c}L_{\text{KL}(\mathbf{c})}+\gamma_{z}L_{\text{KL}(\mathbf{z})}.$ (2)
-We observe that by deactivating the KL term associated with any one of the factors, the likelihood tends to be optimized through the other latent factor.
+We observe that by deactivating the KL term associated with any one of the codes, the likelihood tends to be optimized through the other latent code.
 For instance, if $\gamma_{c}$ is set as 00 and $\gamma_{z}$ has a large value, then the $L_{\text{rec}}$ is optimized through $\mathbf{c}$ only.
 As a consequence, $\mathbf{c}$ captures all the information in $\mathbf{x}$, whereas $\mathbf{z}$ becomes totally independent of the data.
-In such a scenario, $\mathbf{z}$ will be a random factor with no relation to the transformation factor of the data, which is not desirable.
-3.4 Creating Contrastive Pair with respect to Content Latent Factor
-We adopt a contrastive strategy for content-transformation disentanglement.
-We first develop a strategy to impose a structure to the content latent factor with contrastive learning.
+In such a scenario, $\mathbf{z}$ will be a random code with no relation to the transformation of the data, which is not desirable.
+3.4 Creating Contrastive Pair with respect to Content
+We adopt a contrastive strategy for c-z disentanglement.
+We first develop a strategy to impose a structure to the content latent code with contrastive learning.
 To this end, we create negative and positive pairs for each data sample w.r.t.
-the content factor.
-Hence, we randomly pick two samples $\mathbf{x}$ and $\mathbf{x}_{\text{neg}(\mathbf{c})}$, where $\mathbf{x},\mathbf{x}_{\text{neg}(\mathbf{c})}\in\mathcal{X}$ and regard them as negative pairs of each other w.r.t the content factor.
+the content code.
+Hence, we randomly pick two samples $\mathbf{x}$ and $\mathbf{x}_{\text{neg}(\mathbf{c})}$, where $\mathbf{x},\mathbf{x}_{\text{neg}(\mathbf{c})}\in\mathcal{X}$ and regard them as negative pairs of each other w.r.t the content code.
 Such strategy is commonly used in traditional self-supervised learning approaches [16, 24].
@@
 To generate positive pairs w.r.t.
-the content factor, we pick any sample $\mathbf{x}$ and transform it with a transformation $T\in\mathcal{T}$ to generate $\mathbf{x}_{\text{pos}(\mathbf{c})}$.
-We use the encoder $q_{\psi}$ to generate content factors $\mathbf{c}$, $\mathbf{c}_{\text{pos}}$, and $\mathbf{c}_{\text{neg}}$ respectively from $\mathbf{x}$, $\mathbf{x}_{\text{pos}(\mathbf{c})}$, and $\mathbf{x}_{\text{neg}(\mathbf{c})}$.
+the content code, we pick any sample $\mathbf{x}$ and transform it with a transformation $T\in\mathcal{T}$ to generate $\mathbf{x}_{\text{pos}(\mathbf{c})}$.
+We use the encoder $q_{\psi}$ to generate content codes $\mathbf{c}$, $\mathbf{c}_{\text{pos}}$, and $\mathbf{c}_{\text{neg}}$ respectively from $\mathbf{x}$, $\mathbf{x}_{\text{pos}(\mathbf{c})}$, and $\mathbf{x}_{\text{neg}(\mathbf{c})}$.
 We use a contrastive loss to penalize the distance between $\mathbf{c}$ and $\mathbf{c}_{\text{pos}}$ and the similarity between $\mathbf{c}$ and $\mathbf{c}_{\text{neg}}$.
-Our contrastive loss for content latent factor can be written as:
-$L_{\text{con}(\mathbf{c})}=L_{\text{dist}}(\mathbf{c},\mathbf{c}_{\text{pos}})+L_{\text{sim}}(\mathbf{c},\mathbf{c}_{\text{neg}})$ (3)
+Our contrastive loss for content latent code can be written as: $L_{\text{con}(\mathbf{c})}=L_{\text{dist}}(\mathbf{c},\mathbf{c}_{\text{pos}})+L_{\text{sim}}(\mathbf{c},\mathbf{c}_{\text{neg}})$.
 To implement $L_{\text{dist}}$ and $L_{\text{sim}}$, we use mean absolute cosine distance and mean absolute cosine similarity, respectively.
-3.5 Creating Contrastive Pair with respect to Transformation Latent Factor
-Imposing contrastive loss on the content latent factor satisfies condition 1 of the content-transformation disentanglement (Section 3.1).
-However, the content-transformation disentanglement remains only partial since it does not satisfy condition 2.
+3.5 Creating Contrastive Pair with respect to Transformation
+Imposing contrastive loss on the content code satisfies condition 1 of the c-z disentanglement (Section 3.1).
+However, the c-z disentanglement remains only partial since it does not satisfy condition 2.
 To solve this issue, we design another contrastive loss w.r.t.
-the transformation latent factor.
-Designing a contrastive loss that explicitly enforces condition 2 is not possible, given the transformation that may equalize the transformation factor of two samples can not be approximated.
+the transformation code.
+Designing a contrastive loss that explicitly enforces condition 2 is not possible, given the transformation code that may equalize the transformation of two samples can not be approximated.
 To this end, we design a contrastive loss that implicitly encourages condition 2.
@@
 The negative pair of a sample $\mathbf{x}\in\mathcal{X}$ w.r.t.
-transformation latent factor is generated while creating a positive pair of it w.r.t content factor.
+transformation is generated while creating a positive pair of it w.r.t content.
 This is because a transformation $T$ on $\mathbf{x}$ alters its transformation, but not its contents.
@@
 transformation is very challenging as they can not be created in the input space, unlike others.
-To this end, we adopted a strategy of creating pairs of samples from the same distribution in the transformation factor embedding space.
+To this end, we adopted a strategy of creating pairs of samples from the same distribution in the transformation embedding space.
 We randomly sample $\mathbf{z}^{(1)}$ and $\mathbf{z}^{(2)}$ from $\mathcal{N}(0,1)$ and use the decoder $p_{\theta}$ to generate samples $\mathbf{x}^{(1)}_{\text{pos}(\mathbf{z})}$ and $\mathbf{x}^{(2)}_{\text{pos}(\mathbf{z})}$ from them (Fig.
 2 (c)).
-In the decoder, as content input, we use content factor $\mathbf{c}^{(1)}$ and $\mathbf{c}^{(2)}$ obtained by encoding two random samples $\mathbf{x}^{(1)}$ and $\mathbf{x}^{(2)}$.
+In the decoder, as content input, we use content codes $\mathbf{c}^{(1)}$ and $\mathbf{c}^{(2)}$ obtained by encoding two random samples $\mathbf{x}^{(1)}$ and $\mathbf{x}^{(2)}$.
 Consequently, $\mathbf{x}^{(1)}_{\text{pos}(\mathbf{z})}=p_{\theta}(\mathbf{c}^{(1)},\mathbf{z}^{(1)})$ and $\mathbf{x}^{(2)}_{\text{pos}(\mathbf{z})}=p_{\theta}(\mathbf{c}^{(2)},\mathbf{z}^{(2)})$ serve as the positive pairs w.r.t.
-transformation latent factor.
+transformation.
 We then use the encoder $q_{\psi}$ to generate $\mathbf{z}$, $\mathbf{z}_{\text{neg}}$, $\mathbf{z}^{(1)}_{\text{pos}}$, and $\mathbf{z}^{(2)}_{\text{pos}}$ from $\mathbf{x}$, $\mathbf{x}_{\text{neg}(\mathbf{z})}$, $\mathbf{x}^{(1)}_{\text{pos}(\mathbf{z})}$, and $\mathbf{x}^{(2)}_{\text{pos}(\mathbf{z})}$ respectively.
 We use a contrastive loss to penalize the distance between $\mathbf{z}^{(1)}_{\text{pos}}$ and $\mathbf{z}^{(2)}_{\text{pos}}$ and similarity between $\mathbf{z}$ and $\mathbf{z}_{\text{neg}}$.
-Our contrastive loss for transformation latent factor can be written as:
-$L_{\text{con(z)}}=L_{\text{dist}}(\mathbf{z}^{(1)}_{\text{pos}},\mathbf{z}^{(2)}_{\text{pos}})+L_{\text{sim}}(\mathbf{z},\mathbf{z}_{\text{neg}})$ (4)
+Our contrastive loss for transformation can be written as: $L_{\text{con(z)}}=L_{\text{dist}}(\mathbf{z}^{(1)}_{\text{pos}},\mathbf{z}^{(2)}_{\text{pos}})+L_{\text{sim}}(\mathbf{z},\mathbf{z}_{\text{neg}})$.
 Similar to $L_{\text{con}(\mathbf{c})}$, we use mean absolute cosine distance and mean absolute cosine similarity to implement $L_{\text{dist}}$ and $L_{\text{sim}}$, respectively.
+Why rotation?: We used rotation as a transformation to create contrastive pairs because only rotation provided generalization to other shape-based transformations in the dataset, compared to other transformations such as translation, scale, blur, or color-based transformations (Details in Appendix A.2.3).
 Objective Function: In summary, we train the encoder and decoder networks by simultaneously minimizing the loss components. Our overall objective function to minimize is:
 $L=L_{\text{vae}}+L_{\text{con(c)}}+L_{\text{con(z)}}$ (3)
+We optimize $L_{\text{vae}}$ for both $\mathbf{x}$ and the transformed image $\mathbf{x}_{\text{pos}(\mathbf{c})}$ (same as $\mathbf{x}_{\text{neg}(\mathbf{z})}$), which yielded better experimental result.
 4 Experiments & Results
-Table 1: Quantitative Results of unsupervised Content-Transformation Disentangling Methods on MNIST and LineMod datasets.
-This evaluation only shows the disentanglement performance w.r.t ground truth content factors.
-We report mean $\pm$ std.
+Table 1: Quantitative Results of unsupervised Content-Transformation Disentangling Methods.
+Among these methods, SpatialVAE and Harmony put constraints on the $\mathbf{z}$ code dimension given their explicit parameterization, others do not.
+The std.
 dev.
-over model training by setting 3 different random seeds.
-Here, for Harmony and SpatialVAE, the transformation latent factor is restricted to dimension 3.
-For the other methods, there is no such restriction on the transformation latent factor.
-For SimCLR, (Disc) means discriminative model (without any reconstruction), and (Gen) means Generative model (with reconstruction).
-MNIST LineMod
-Method $D_{\text{score}}(c|c)(\uparrow)$ $D_{\text{score}}(c|s)(\downarrow)$ $D_{\text{score}}(c|c)(\uparrow)$ $D_{\text{score}}(c|s)(\downarrow)$
-SpatialVAE [2] 0.81 $\pm$ 0.01 0.28 $\pm$ 0.02 0.95 $\pm$ 0.02 0.32 $\pm$ 0.01
-Harmony [4] 0.82 $\pm$ 0.01 0.31 $\pm$ 0.02 0.90 $\pm$ 0.01 0.56 $\pm$ 0.02
-SimCLR (Disc) with rotation [15] 0.58 $\pm$ 0.01 0.60 $\pm$ 0.01 0.62 $\pm$ 0.01 0.40 $\pm$ 0.01
-SimCLR (Gen) with rotation [15] 0.53 $\pm$ 0.01 0.67 $\pm$ 0.02 0.61 $\pm$ 0.02 0.79 $\pm$ 0.01
-VAE with two latent space 0.63 $\pm$ 0.00 0.63 $\pm$ 0.00 0.87 $\pm$ 0.00 0.87 $\pm$ 0.00
-VITAE [1] 0.77 $\pm$ 0.00 0.32 $\pm$ 0.00 0.92 $\pm$ 0.00 0.90 $\pm$ 0.00
-DualContrast (w/o $L_{\text{con(c)}}$) 0.87 $\pm$ 0.02 0.21 $\pm$ 0.02 0.86 $\pm$ 0.01 0.31 $\pm$ 0.02
-DualContrast (w/o $L_{\text{con(z)}}$) 0.79 $\pm$ 0.04 0.85 $\pm$ 0.02 0.79 $\pm$ 0.03 0.86 $\pm$ 0.03
-DualContrast 0.89 $\pm$ 0.01 0.31 $\pm$ 0.02 0.95 $\pm$ 0.01 0.48 $\pm$ 0.02
-Benchmark Datasets: We report results on MNIST [34], LineMod [35], Starmen [36], and a realistically simulated protein subtomogram dataset.
-Only LineMod features RGB images among these datasets, while the other three consist of grayscale images.
+over model training by setting 3 different random seeds remains within ±$\pm 0.04$.
+$D_{\text{score}}$ is written as $D$.
+For $D(c|c)$ and $D(z|z)$, higher is better.
+For $D(c|z)$ and $D(z|c)$, lower is better.
+For $SAP(c)$ and $SAP(z)$, higher is better.
+MNIST LineMod Protein Subtomogram
+Method $D(c|c)$ $D(c|z)$ $SAP(c)$ $D(c|c)$ $D(c|z)$ $SAP(c)$ $D(c|c)$ $D(c|z)$ $SAP(c)$ $D(z|z)$ $D(z|c)$ $SAP(z)$
+SpatialVAE 0.81 0.28 0.53 0.95 0.32 0.63 0.69 0.93 0.24 0.71 0.57 0.14
+Harmony 0.82 0.31 0.51 0.90 0.56 0.34 0.95 0.01 0.94 0.52 0.90 0.38
+SimCLR (Discriminative) 0.58 0.60 0.02 0.62 0.40 0.22 0.39 0.69 0.30 0.51 0.52 0.01
+SimCLR (Generative) 0.53 0.67 0.14 0.61 0.79 0.18 0.53 0.59 0.06 0.55 0.54 0.01
+VAE with $2$ latent space 0.63 0.63 0.0 0.87 0.87 0.0 0.79 0.82 0.03 0.85 0.85 0.0
+VITAE 0.77 0.32 0.45 0.92 0.90 0.02 - - - - - -
+DualContrast (w/o $L_{\text{con(c)}}$) 0.87 0.21 0.66 0.86 0.31 0.55 0.93 0.81 0.13 0.78 0.81 0.03
+DualContrast (w/o $L_{\text{con(z)}}$) 0.79 0.85 0.06 0.79 0.86 0.07 0.99 0.86 0.13 0.86 0.63 0.23
+DualContrast 0.89 0.31 0.58 0.95 0.48 0.47 1.00 0.56 0.44 0.86 0.64 0.22
+Benchmark Datasets: We report results on MNIST [25], LineMod [26], and a realistically simulated protein subtomogram dataset.
+Only LineMod features RGB images among these datasets, while the others consist of grayscale images.
 The subtomogram is a 3D volumetric dataset, whereas the others are 2D images.
-Baselines: We used the unsupervised content-transformation disentanglement methods, SpatialVAE [2], Harmony [4], VITAE [1] with explicit parameterization and InfoNCE loss based standard self-supervised learning method with rotation augmentation [15] as baseline approaches.
+For a detailed discussion of the datasets, we refer to the Appendix.
+Baselines: We used the unsupervised c-z disentanglement methods, SpatialVAE [2], Harmony [4], VITAE [1] with explicit parameterization and InfoNCE loss based standard self-supervised learning method with rotation augmentation [15] as baseline approaches.
 Several generic disentangled representation learning methods, such as beta-VAE, Factor-VAE, beta-TCVAE, DIP-VAE, etc., are available that do not aim to disentangle specific factors such as content or transformation in the data but rather aim to disentangle all factors of variation.
@@
 Implementation Details: We implemented the encoder using a Convolutional Neural Network (CNN) and the decoder using a Fully Connected Network (FCN).
-Our experiments used the same latent dimension for content and transformation factors, except for Harmony [4] and SpatialVAE [2], where a specific dimension needs to be used for transformation latent factors.
+Our experiments used the same latent dimension for content and transformation codes, except for Harmony [4] and SpatialVAE [2], where a specific dimension needs to be used for transformation codes.
 We train our models for $200$ epochs with a learning rate of $0.0001$ in NVIDIA RTX A500 GPUs and AMD Raedon GPUs.
 We optimize the model parameters with Adam optimizer.
-Evalutation Metrics: We assess DualContrast based on two criteria: (1) the quantitative disentanglement of content and transformation factors and (2) the qualitative capability of image generation through content-transformation transfer.
-Figure 2: Qualitative content-transformation transfer based image generation results for MNIST.
-(a), (b), (c), and (d) shows the results with Harmony [4], Spatial-VAE [2], VITAE [1], and DualContrast respectively.
+The hyperparameters $\gamma_{z}$ and $\gamma_{c}$ controls how close the z and c distribution will be to $\mathcal{N}(0,I)$.
+We found setting a small value $\approx 0.01$ for both provides optimal results in our experiments.
+Further details of the overall implementation are provided in the Appendix A.3.
+Evalutation: We evaluate the methods based on two criteria: (1) the informativeness of the predicted $\mathbf{c}$ and $\mathbf{z}$ codes w.r.t the ground truth $\mathbf{c}$ and $\mathbf{z}$ factors, (2) the separateness of of the predicted $\mathbf{c}$ and $\mathbf{z}$ codes.
+We performed these evaluations both quantitatively and qualitatively.
+For quantitative evaluations, there exists several metrics, e.g.
+, MIG score, $D_{\text{score}}$, SAP score, etc.
+[27] demonstrated that these metrics are highly correlated.
+To this end, we only used $D_{\text{score}}$ and SAP score to measure the disentanglement.
+$D_{\text{score}}$ is simply a measurement of how well the ground truth (gt) factors can be predicted from the corresponding latent codes.
+In our scenario, there are four such quantities - (a) predictivity of content gt given $\mathbf{c}$ code $D_{\text{score}}(c|c)$ or $\mathbf{z}$ code $D_{\text{score}}(c|z)$ (b) predictivity of transformation gt given $\mathbf{c}$ code $D_{\text{score}}(z|c)$ or $\mathbf{z}$ code $D_{\text{score}}(z|z)$.
+For MNIST and LineMod, we do not have any transformation gt, so we only reported values for (a).
+Similar to [15], we use linear Logistic Regression Classifiers to measure predictivity ($D_{\text{score}}$).
+The higher the $D_{\text{score}}$, the more informative is the latent code w.r.t.
+the gt.
+SAP score for a gt factor is defined as the difference between the highest and second highest predictivities for it given any latent codes.
+In our case, SAP(c) can be defined as $|D_{\text{score}}(c|c)-D_{\text{score}}(c|z)|$ and SAP(z) as $|D_{\text{score}}(z|z)-D_{\text{score}}(z|c)|$.
+The higher the SAP score, higher the separateness of the latent codes.
+For qualitative evaluations, we performed downstream tasks like generating images through unsupervised content-transformer transfer and latent space visualization.
+For protein subtomograms, we performed subtomogram averaging [28] to identify proteins resulting from the latent space clusters.
+4.1 DualContrast disentangles several writing styles from MNIST images
+Figure 3: Qualitative Results of Unsupervised c-z Disentanglement on MNIST obtained by (a) Harmony, (b) SpatialVAE, (c) VITAE, and (d) DualContrast, respectively.
+Images are generated by the Decoders given content (c) code from the leftmost column images and transformation (z) code from the topmost row images.
+Similar to the related works, we start our experiments with the MNIST dataset of handwritten digits.
+We performed a quantitative comparison between DualContrast and the baseline approaches, as shown in Table 1.
+We observed that DualContrast achieves higher $D_{score}(c|c)$ compared to other explicitly parameterized approaches [4, 2, 1].
+However, since transformation labels are not present for MNIST, relying entirely on the content prediction performance for disentanglement might be misleading.
+To this end, we report qualitative results for baseline methods and our approach DualContrast (Fig.
+3 and Appendix Fig.
+10).
+In generating images with varying c and z codes, we observe that Harmony, SpatialVAE, and C-VITAE generate many images with erroneous content and transformations.
+On the other hand, DualContrast does not make such mistakes.
+This again suggests better c-z disentanglement with our approach than existing explicit parameterization approaches.
+Moreover, DualContrast disentangles more than in-plane rotation for MNIST.
+For the digit 00 marked with green box in Fig.
+3, Harmony and SpatialVAE simply rotate the image, not capturing the actual writing style of the digit $2$ in the top row.
+VITAE somewhat represented the transformation with its explicit modeling of piecewise linear transformation.
+On the other hand, DualContrast properly captured the writing style of above digit $2$.
+We included additional qualitative results, including the $\mathbf{c}$ and $\mathbf{z}$ latent space visualization, in the Appendix A.4.
+4.2 DualContrast Disentangles ViewPoint from Linemod Object Dataset
+Figure 4: Qualitative Results of Unsupervised c-z Disentanglement on LineMod obtained by (a) Harmony, (b) SpatialVAE, (c) VITAE, and (d) DualContrast, respectively.
+Images are generated by the Decoders given c code from the leftmost column images and z code from the topmost row images.
+Additional content-transformation transfer results are available on Appendix Fig.
+12.
+LineMod is an RGB object recognition and 3D pose estimation dataset [26] with 15 unique object types.
+The objects are segmented from real-world scenes, and the segmented images visualize the objects from different viewpoints.
+The entire dataset has $,31$ 313 images per object category.
+We used $1,000$ images per category for training and the remaining for testing.
+We evaluated the scores $D_{\text{score}}$ (Table 1) for DualContrast and baseline approaches, where DualContrast clearly shows the best performance.
+During evaluation, we used the object identities as ground truth content factors.
+We also performed qualitative measurement similar to MNIST and provided the results in Fig.
+4 and Appendix 12.
+We observe that VITAE distorts the images significantly while performing image generation with c-z transfer.
+On the other hand, DualContrast does not face such issues.
+SpatialVAE and Harmony only disentangles in-plane rotation given their design; so changing the z code in these methods only rotates the content sample.
+However, DualContrast can disentangle different viewpoints of the objects as the z code and changing z changes the viewpoint of objects- reflecting the actual transformation present in the dataset.
+4.3 DualContrast disentangles protein composition from conformations in Cryo-ET subtomograms and enables their precise identification
+We created a realistically simulated [29] cellular cryo-ET subtomogram dataset of 18,000 samples belonging to 3 different protein identities - Nucleosome, Sars-Cov-2 spike protein, and Fatty Acid Synthase (FAS) unit.
+We collected 6 different conformations or compositional states for each protein identity from RCSB PDB [30].
+Among them, only nucleosome composition has subtly visible differences across all the 6 states.
+For each conformation per protein identity, 1,000 subtomograms were generated.
+Each subtomogram is of size $32^{3}$ and contains a protein in random orientation and shift with a very low signal-to-noise ratio (Fig.
+5).
+Further details on the dataset properties are provided in the Appendix A.4.
+Figure 5: Disentanglement of composition and conformations in cellular subtomogram dataset with slice-by-slice visualization of $4$ sample subtomograms.
+UMAP embedding of c codes in (a) SpatialVAE, (b) Harmony, and (c) DualContrast.
+(d) Slice-by-slice visualization of x-y slices in $4$ sample subtomograms.
+(e) UMAP embedding of c codes in Harmony trained only nucleosome subtomograms.
+(f) UMAP embedding of z codes in Harmony trained with all subtomograms.
+Figure 6: (a,b,c) Structures obtained with RELION refinement for each (4) cluster of subtomograms, whereas the clustering is performed using Gaussian Mixture Modeling (GMM) on $\mathbf{c}$ codes, predicted by (a) SpatialVAE (b) Harmony, and (c) DualContrast.
+(d, e) Structures obtained with RELION refinement for the nucleosome subtomograms (identified with previous step) using GMM on (d) Harmony content ($\mathbf{c}$) codes, and (e) DualContrast transformation ($\mathbf{z}$) codes.
+We trained DualContrast and baseline approaches (except VITAE since it could not be trained without designing a 3D CPAB transformation) against the protein subtomogram dataset.
+We provided the quantitative results in Table 1 and qualitative results in Fig.
+5 and Fig.
+6 respectively.
+For qualitative results, we first perform UMAP visualization of the latent codes obtained by the models.
+We observed that the UMAP of c codes in DualContrast perfectly clustered the protein identities from the dataset, profoundly disentangling the transformations (Fig.
+5(c)).
+We do not see such clustering for Harmony (Fig.
+5(b)); rather, protein identities with very different compositions get mixed up due to their conformational variation.
+SpatialVAE latent code UMAP could not cluster the proteins at all (Fig.
+5(a)).
+We further observe that DualContrast z code UMAP clusters have similar nucleosome conformations, showing disentanglement of conformation from protein identity.
+On the other hand, in Harmony, the transformation factor only represents rotation and translation by design and can not capture nucleosome conformations.
+Moreover, even its $\mathbf{c}$ code for manually selected (not through automatic clustering) nucleosome subtomograms can not show good clustering of conformations (Fig.
+5(b)).
+To further demonstrate the application of such disentanglement, we perform downstream subtomogram averaging to identify the structures from latent space clusters.
+Subtomogram averaging obtains readily identifiable high SNR structures from multiple low SNR subtomograms through iterative alignment and averaging [28].
+In our experiments, Gaussian Mixture Models (GMM) were applied to content ($\mathbf{c}$) codes predicted by models, generating clusters of subtomograms that were averaged using RELION [31] (Fig.
+6).
+SpatialVAE failed to identify proteins, while Harmony identified the FAS, nucleosome, and $2$ spike proteins, though $1$ spike protein appeared as an unrealistic mixture of FAS and spike protein.
+DualContrast successfully identified the FAS, nucleosome, and $2$ distinct spike protein compositions.
+For nucleosome conformations, GMM clustering of Harmony’s $\mathbf{c}$ codes for nucleosome cluster subtomograms revealed only $1$ nucleosome conformation but mistakenly included a spike protein.
+In contrast, clustering DualContrast’s $\mathbf{z}$ codes followed by RELION averaging identified $4$ distinct nucleosome conformations with subtle structural changes, showcasing its ability to disentangle protein composition and conformation, which is unattained by the other existing methods.
+Figure 7: Content-transformation transfer results from ablation analysis.
+(a) and (b) shows the results when the model is trained with $L_{\text{VAE}}+L_{\text{con (z)}}$ and $L_{\text{VAE}}+L_{\text{con (c)}}$ respectively.
+Ablation Study: To evaluate the individual contribution of the contrastive losses, we conduct both quantitative and qualitative ablation analyses of DualContrast.
+We trained (1) DualContrast without any contrastive loss, which is basically a VAE with two latent spaces, (2) DualContrast with only $L=L_{\text{VAE}}+L_{\text{con (z)}}$, and (3) $L=L_{\text{VAE}}+L_{\text{con (c)}}$.
+We qualitatively and quantitatively evaluated each model.
+We show the quantitative results in Table 1 and qualitative results of MNIST in Fig.
+7.
+We provide qualitative results and a detailed discussion on the ablation in the Appendix A.4.4.
+5 Discussions & Limitations
+We introduced an unsupervised content-transformation disentanglement method that, for the first time, does not rely on labels or explicit parameterization of transformations.
+Our method successfully disentangled transformations that cause subtle pixel-space changes, such as variations in writing styles (MNIST), viewpoint changes (LineMod), and, most importantly, conformational changes in proteins from protein-mixture cryo-ET subtomogram datasets.
+However, it is not guaranteed to disentangle all transformations, particularly those causing large changes in the pixel space.
+Instead, the method may classify transformations causing large pixel-space changes as content.
+Nonetheless, this aligns with scientific imaging, where large changes often reflect compositional shifts and subtle changes represent conformational variations.
+Disentangling transformations with large pixel-space changes in the pixel space without any explicit parameterization is extremely challenging and might not be practically achievable, but it remains an avenue for future research.
+6 Conclusion
+This work focuses on a challenging setting of the unsupervised content-transformation disentanglement problem in scientifically important shape focused image datasets, where the transformation latent code is not explicitly parameterized.
+To tackle this problem, we propose a novel method termed DualContrast.
+DualContrast employs generative modeling with a novel contrastive learning strategy that creates positive and negative pairs for content and transformation latent codes.
+Our extensive experiments show DualContrast’s effective disentanglement of challenging transformations across various shape-based image datasets; including simulated cellular subtomograms, where it solved the unexplored problem of isolation of protein conformations from protein identities, as proof of principle.
+7 Acknowledgement
+This work was supported in part by U.S.
+NIH grants R01GM134020 and P41GM103712, NSF grants DBI-1949629, DBI-2238093, IIS-2007595, IIS-2211597, and MCB-2205148.
+This work was supported in part by Oracle Cloud credits and related resources provided by Oracle for Research, and the computational resources support from AMD HPC Fund.
+MRU were supported in part by a fellowship from CMU CMLH.
+Appendix A Appendix
+Overview
+  • Appendix A.1 contains additional discussion on related works.
+  • Appendix A.2 contains an additional mathematical explanation of DualContrast and Baseline methods.
+  • Appendix A.3 contains additional details on the experiments.
+  • Appendix A.4 contains additional results and description on the datasets.
+A.1 Detailed Related Works
+Disentangled Representation Learning: PCA [32] and ICA [33] can be regarded as very preliminary work in the domain of disentangled representation learning.
+However, these methods assume linear subspace and do not work well for non-linear high-dimensional datasets.
+Deep learning-based approaches like Info-GAN [34], $\beta$-VAE [35], and their variants [36, 37, 38, 39] have recently gained wide attention as generic approaches for learning disentangled representations.
+Most of these works manipulated the variational bottleneck to achieve disentanglement of the latent codes.
+However, these works do not aim toward disentangling any specific factor, e.g.
+, content, group, style, transformation, etc., from the latent codes.
+Instead, they generate a series of images by traversing through each dimension of the latent space while keeping the remaining dimensions fixed.
+Thus, they infer the semantic meaning of each dimension of the learned latent factor.
+Consequently, these methods do not perform well in disentangling any specific generative factor compared to those that aim to disentangle several (two in most cases) specific generative factors [4, 1, 2].
+Unlike these methods, our method specifically disentangles the content and transformation factors of data samples, whereas the content and transformation are defined in Section 3.1.
+cryo-EM Heterogeneous Reconstruction: There exists several works on single particle cryo-EM and cryo-ET reconstruction, e.g.
+, cryoDRGN2 [40], cryoFIRE [3], cryoAI [41], etc.
+that performs amortized inference of transformation ($SO(3)\times d^{2}$) and latent space representing content.
+However, these works mainly focus on 2D-to-3D reconstruction instead of content-transformation disentanglement.
+Our work, on the contrary, focuses on content-transformation disentanglement.
+Though we use reconstruction loss to maximize informativeness of content and transformation factors, our reconstruction is 2D-to-2D or 3D-to-3D, unlike the aforementioned works.
+Also, our transformation factor is implicit and not explicitly limited to $SO(3)\times d^{2}$.
+Shape Analysis: Disentangling content and transformation latent factors have special significance in the domain of shape analysis.
+Consequently, shape representation learning, modeling, and analysis [42, 43, 44, 5, 45, 12] are closely related to our work.
+Even for shape analysis, PCA can be regarded as one of the primitive methods.
+Even now, PCA is widely used in the shape analysis of protein complexes [46].
+Recently, Huang et al.
+[45] demonstrated that PCA with two components on the latent factor learned by an auto-encoder corresponds to content (shape) and style (pose) in 3D human mesh datasets.
+Nevertheless, PCA is a linear method assuming linear subspaces, which often does not hold true.
+A line of shape analysis research [12, 47, 43, 5] has been performed for non-linear disentanglement of content and style factors in 3D mesh or point cloud datasets.
+The goal of these works is to reduce per-vertex reconstruction loss of 3D meshes or point clouds for content-style-specific generation.
+These works claim unsupervised disentanglement as they do not require ground truth factors.
+However, they use the identity information of meshes, which is directly associated with content code.
+In contrast, our method does not require identity information apriori to learn latent codes specific to shape and code.
+Moreover, the mentioned works specifically investigate mesh-specific geodesic losses to achieve minimal per-vertex mesh reconstruction.
+On the other hand, we do not specifically aim to design mesh-specific losses in this work, rather, we propose a generic content-transformation disentanglement approach that can be applied to 3D mesh datasets with necessary modification in the model architecture.
+A.2 Method
+A.2.1 Content-Transformation disentanglement with variational autoencoders (VAE)
+A standard Variational Autoencoder (VAE) presumes data $\mathbf{x}$ to be generated by latent variable $\mathbf{z}$, whereas a standard Gaussian prior is assumed for $\mathbf{z}$.
+[…]
+$p(\mathbf{z})$ $=\mathcal{N}(0,\mathbf{I}_{d})$
+We extended the standard VAE to a two-latent variable setting. We assume latent variables $\mathbf{z}$ and $\mathbf{c}$ to generate the data $\mathbf{x}$.
+[…]
+This setting is similar to VITAE [1], SpatialVAE [2], and Harmony [4].
+However, in SpatialVAE [2] and Harmony [4], latent factor $\mathbf{z}$ is restricted as rotation and parameterized transformations, respectively.
+In SpatialVAE,
+$p(\mathbf{z})$ $=\text{Unif}(a,b)$ (4)
+$\theta$ $\sim p(\mathbf{z}|\mathbf{x})$ (5)
+$\mathbf{x}_{\text{cord}}$ $=\mathcal{R}([-1,1]^{d\times d};\theta)$ (6)
+[…]
+where $a$ and $b$ are specified constants, $\theta$ are transformation (2D rotation and translation) parameters, $\mathcal{R}$ is the corresponding transformation operator.
+On the other hand, in Harmony,
+$\theta$ $=\mathbb{I}(\mathbf{z}|\mathbf{x})$
+$\mathbf{x}^{\prime}$ $=\mathcal{T}(x;\theta)$
+[…]
+where $\mathbb{I}$ is an identity function, $\theta$ are transformation parameters and $\mathcal{T}$ is the corresponding transformation operator.
+Unlike these two methods, in VITAE [1] and our proposed DualContrast, we use standard Gaussian priors for latent codes $\mathbf{z}$ and $\mathbf{c}$.
+$p(\mathbf{z})$ $=\mathcal{N}(0,\mathbf{I}_{d})$
+$p(\mathbf{c})$ $=\mathcal{N}(0,\mathbf{I}_{d})$
+However, in VITAE [1], $\mathbf{z}$ is used to explicitly sample continuous piecewise affine velocity (CPAB) transformation parameter $\theta$, and $\mathbf{c}$ is used to sample appearance samples $\mathbf{x}^{\prime}$.
+$\theta$ $\sim p(\mathbf{x}|\mathbf{z})$
+$\mathbf{x}^{\prime}$ $\sim p(\mathbf{x}|\mathbf{c})$
+$\mathbf{x}$ $=\mathcal{T}(\mathbf{x}^{\prime};\theta)$
+where $\mathcal{T}$ is the transformation operator for CPAB transformation. CPAB transformation parameter is highly expressive compared to affine transformation parameters used in spatialVAE.
+Contrary to VITAE [1], we do not use $\mathbf{z}$ to sample any transformation parameters explicitly; rather use both $\mathbf{z}$ and $\mathbf{c}$ to generate $\mathbf{x}$.
+To this end, we use a contrastive learning strategy that is described in Section 3 of the main paper.
+Without explicitly sampling any transformation parameter, we improve the expressiveness of our transformation latent factor $\mathbf{z}$ even more than the CPAB transformation used in VITAE [1].
+A.2.2 Feature Suppression of SimCLR and MoCo contrastive losses:
+The contrastive losses used in popular self-supervised learning methods SimCLR [16] or MoCo [24] as did not help much in disentangling content and transformation in our experiments.
+It has been demonstrated that these methods often learn nuisance image features or noise to obtain a shortcut solution to the contrastive objective [8].
+This phenomenon is referred to as feature suppression of contrastive objectives.
+We found that using reconstructive loss was necessary to prevent the feature suppression problem.
+A.2.3 Choice of Transformation to create Contrastive pairs
+Figure 8: Content-transformation transfer based image generation results using different transformation functions ($T(x)$) to create contrastive pairs.
+Only rotation ensures transformation factor $z$ to capture object viewpoints- the transformation factor present in the original dataset.
+We leveraged different transformation functions $T(x)$ to create contrastive pairs in DualContrast for LineMod RGB object dataset.
+We used rotation, translation, scaling, blur, saturation, and hue as $T(x)$.
+We performed both qualitative (Table 2) and quantitative analysis on the effect of different $T(x)$ for content-transformation disentanglement in DualContrast.
+We observe that using Scale or Blur makes the transformation factor z uninformative of the data and it does not capture anything at all.
+Consequently, changing this z factor while generating images does not change the image at all for these two codes (Figure 8).
+On the other hand, using translation shows small negligible differences in the c-z transfer-based image generation.
+Color-based transformations like saturation and Hue only change the color of the generated image, instead of affecting its shape-based transformation.
+Only rotation provides generalization of z and enables z to capture viewpoint transformations present in the dataset.
+Table 2: Transformation Factors and Corresponding $D_{\text{score}}(c|z)$ and $D_{\text{score}}(c|c)$ values.
+Transformation Factor $D_{\text{score}}(c|z)$ $D_{\text{score}}(c|c)$
+Rotation 0.48 0.95
+Translation 0.65 0.98
+Scale 0.52 0.91
+Contrast 0.51 0.93
+Saturation 0.71 0.86
+Hue 0.61 0.88
+Blur 0.47 0.92
+A.3 Experiments
+A.3.1 Implementation Details
+Figure 9: Visualization of the creation of contrastive pairs for MNIST.
+(a) Creating positive pair with respect to (w.r.t) content factor and negative pair w.r.t.
+transformation.
+(b) Creating negative pair w.r.t.
+content.
+(c) Creating positive pair w.r.t.
+transformation.
+We show the processes for a batch of MNIST digits with a batch size of 5.
+We implemented our model in Pytorch (version 1.9.0).
+We used a convolutional neural network (CNN) (3 convolutional layers for MNIST, 4 for others) to implement the encoder and a fully connected network (FCN) (5 layers) to implement the decoder.
+For subtomograms, we used a 3D convolutional network for the encoder.
+We do not use any pooling layers in our networks.
+While training the models, we use a batch size of 100 and an Adam optimizer with a learning rate of $0.0001$.
+We used a linear learning rate scheduler that decays the learning rate of each parameter group by 0.1 every 50 epochs.
+We trained our models for 200 epochs.
+We used the same setting for our models and the baseline models.
+We used NVIDIA RTX A500 and AMD Radeon GPUs to train the models.
+Choice of latent dimension: For Harmony [4] and SpatialVAE [2], the transformation latent factor is restricted to dimension $3$.
+For VITAE [1], SimCLR [15], and our DualContrast, there is no such restriction on the transformation latent factor dimension and same dimension was used as the content factor.
+For all the methods, the dimension of the content factor was set as $10$ for MNIST, $50$ for subtomogram dataset.
+For hyperparameters $\gamma_{\mathbf{c}}$ and $\gamma_{\mathbf{z}}$, we set a small value ($\approx 0.01$) in our experiments.
+The values determine how strictly the content factor and the transformation factor should mimic the prior standard multivariate gaussian distribution.
+A.4 Additional Results
+A.4.1 MNIST
+We use the commonly used MNIST dataset to initialize our experiments.
+MNIST is a dataset in the public domain that the research community has extensively used.
+It contains grayscale images of handwritten digits.
+Each image is of size $28\times 28$.
+The training set contains 60,000 images, whereas the test set contains 10,000.
+We use the same train test split for our experiments.
+We show a full visualization of content-transformation transfer based image generation of Figure 3 in Figure 10.
+We also include tSNE embedding of the content codes inferred by the models on the MNIST test dataset associated with class labels (Figure 11).
+On the embedding space, DualContrast clearly shows superior clustering performance.
+((a))
+((b))
+((c))
+((d))
+Figure 10: Content-transformation transfer results from ablation analysis (a), (b), (c), and (d) shows the results with Harmony [4], Spatial-VAE [2], VITAE [1], and DualContrast respectively.
+When generating image grids, the transformation factor is uniform across rows, and the content factor is uniform across columns.
 Erroneous generations (both in terms of content and transformation) are marked within red boxes.
-For (1), there exists several metrics, e.g., MIG score, $D_{\text{score}}$, etc.
-Locatello et al [37] demonstrated that these metrics are highly correlated.
-To this end, similar to [1], we only use $D_{\text{score}}$ to measure the disentanglement.
-$D_{\text{score}}$ is simply a measurement of how well the ground truth factors can be predicted from the corresponding latent factors.
-In our scenario, there are four such quantities - (a) predictivity of content labels given content factor $D_{\text{score}}(c|c)$ or transformation factor $D_{\text{score}}(c|s)$ (b) predictivity of transformation labels given content factor $D_{\text{score}}(s|c)$ or transformation factor $D_{\text{score}}(s|s)$.
-For MNIST and LineMod, we do not have any transformation labels, so we only reported values for (a).
-Similar to [15], we use linear Logistic Regression Classifiers to measure predictivity ($D_{\text{score}}$) across all our experiments.
-Figure 3: tSNE embedding plots of content latent factor learned by the unsupervised content-style disentanglement methods.
+Figure 11: tSNE embedding plots of content latent factor learned by the unsupervised content-transformation disentanglement methods.
 (a), (b), (c), and (d) shows the results for Harmony [4], Spatial-VAE [2], C-VITAE [1], and DualContrast respectively.
 Overall, DualContrast shows superior performance.
-The quantitative measurement, particularly when transformation labels are not available, might be misleading.
-To this end, we perform qualitative evaluation by generating images while transferring content and transformation from original images from the dataset.
-We take a set of $K$ images $\mathbf{x}$ and create a $K\times K$ grid of $K^{2}$ images using the different combinations of content and transformation factors from those $K$ images.
-In the grid, we keep the content factor and transformation factor consistent across the columns and the rows respectively.
-In other words, the image at position $(i,j)$ in the grid is generated using the content of image $x_{j}$ and transformation of image $x_{i}$.
-For the protein subtomogram dataset, we also provide qualitative results through embedding visualization (Figure 8).
-4.1 DualContrast disentangles writing style from MNIST handwritten digit images
-Similar to [4, 2, 1], we start our experiments with the MNIST dataset of handwritten digits.
-We performed a quantitative comparison between DualContrast and the baseline approaches, as shown in Table 1.
-We observed that DualContrast achieves higher $D_{score}(c|c)$ compared to other explicitly parameterized approaches [4, 2, 1].
-Figure 4: Qualitative Results of Unsupervised Content-Transformation Disentanglement on LineMod obtained by VITAE[1] (a) and DualContrast (b), respectively.
-However, since transformation labels are not present for MNIST, relying entirely on the content prediction performance for disentanglement might be misleading.
-For example, the degenerate solution achieves even lower $D_{score}(c|s)$ as transformation does not represent anything.
-To this end, we report qualitative results for baseline methods and our approach DualContrast (Fig.
-2).
-In generating images with varying content and styles, we observe that Harmony, SpatialVAE, and C-VITAE generate many images with erroneous content and styles.
-On the other hand, DualContrast produces a few minor errors and no extreme mistakes.
-We also include tSNE embedding of the content factors inferred by the models on the MNIST test dataset associated with class labels (Figure 3).
-On the embedding space, DualContrast clearly shows superior clustering performance.
-All of these suggest better content-transformation disentanglement with our approach than existing explicit parameterization approaches.
-4.2 DualContrast Disentangles ViewPoint from Linemod Object Dataset
-Figure 5: Qualitative Results of Unsupervised Content-Transformation Disentanglement on Starmen (a) and (b) shows the results obtained with VITAE [1] and DualContrast, respectively.
-LineMod is an RGB object recognition and 3D pose estimation dataset [35] with 15 unique object types.
+A.4.2 LineMod
+LineMod [26] dataset is originally designed for object recognition and 6D pose estimation.
 It contains 15 unique objects: ‘ape’, ‘bench vise’, ‘bowl’, ‘cam’, ‘can’, ‘cat’, ‘cup’, ‘driller’, ‘duck’, ‘eggbox’, ‘glue’, ‘hole puncher’, ‘iron’, ‘lamp’ and ‘phone’, photographed in a highly cluttered environment.
@@
 The dataset is publicly available at this url.
-The dataset has $,31$ 313 images per object category.
-We used $1,000$ images per category for training and the remaining for testing.
-We evaluated the disentanglement scores $D_{\text{score}}$ (Table 1) for DualContrast and baseline approaches, where DualContrast clearly shows the best performance.
-We used the object identities as content labels during the evaluation.
-We also performed qualitative measurement with image transfer and depicted the result of DualContrast and its most relevant (in terms of the generality of transformations) VITAE in Fig 4.
-We observe that VITAE distorts the images significantly while performing image generation with content-style transfer.
-On the other hand, DualContrast does not face such issues.
-Moreover, DualContrast can disentangle different viewpoints of the objects as the transformation factor using only 2D rotation to create contrastive pairs in its model.
-4.3 DualContrast disentangles Human Deformation from starmen shapes dataset
-Starmen is a 2D synthetic shape dataset of the deformable human body [36] consisting of $10,000$ shape-focused images, which is publicly available at https://zenodo.org/records/5081988 under CC-BY license.
-Each starmen image is a grayscale image of size $64\times 64$.
-They are generated by deforming five control points (two on arms, two on legs, one on head) on template human shapes.
-We randomly took $6,000$ starmen images to train our models and kept the remaining $4,000$ for testing.
-We report the qualitative results by VITAE [1] and DualContrast in Figure 5.
-From the image generation results with the content-transformation transfer, we can observe that DualContrast disentangles the deformation of starmen shapes as transformation and the identity (based on leg thickness) as content, providing a smooth image generation result.
-On the other hand, VITAE [1] does not perform such meaningful content-transformation disentanglement, generating implausible results.
-4.4 DualContrast Disentangles Protein Identity from states in Cryo-ET Subtomograms
-Figure 6: (a) Isosurface visualization of Nucleosome Density Maps.
-Each density map slightly varies in terms of conformation.
-They are associated with the PDB IDs in the figure.
-(b) 3D Slice by slice visualization of several subtomograms of the corresponding nucleosome density maps.
+The dataset is publicly available under MIT License.
+((a))
+((b))
+((c))
+((d))
+Figure 12: Qualitative results of image generation with content-transformation transfer obtained by (a) Harmony [4], (b) SpatialVAE [2], (c) VITAE [1], and (d) DualContrast respectively.
+Harmony and SpatialVAE perform well in reconstruction.
+but can only perform rotation and translation with its explicitly defined transformations.
+On the other hand, VITAE can comparatively perform better disentanglement with very poor reconstruction results, distorting the images.
+On the other hand, DualContrast provides superior content-transformation transfer with optimal performance in reconstruction.
+This dataset has $,31$ 313 images per object category.
+We used $1,000$ images per category for training and used the remaining for testing.
+For many objects, the object region covers only a tiny part of the original image.
+To this end, we cropped the object region from the original image using the segmentation masks provided with the original dataset.
+After cropping the object regions, we padded $8$ pixels to each side of the cropped image and then reshaped the padded image to the size of $(64,64,3)$.
+Thus, we prepared the training and testing datasets for content-transformation disentanglement in LineMod.
+We used the same dataset and train-test splits for our model and the baseline models.
+The associated processing codes are provided in the supplementary material.
+We trained our proposed DualContrast, VITAE [1], SpatialVAE [2], and Harmony [4] on the LineMod dataset.
+We provide qualitative results of image generation with content-transformation transfer in Fig.
+12 obtained with each method.
+It is noticeable that both Harmony [4] and SpatialVAE [2] have shown good performance when it comes to reconstruction.
+However, these two methods can only perform rotation and translation of the objects with explicitly defined transformations and can not capture complex transformations, e.g., projection, viewpoint change, etc.
+Compared to SpatialVAE and Harmony, VITAE [1] can perform better transformation transfer but performs poor reconstruction.
+Nevertheless, DualContrast stands out for its superior ability to perform transformation transfer while ensuring optimal performance in reconstruction.
+A.4.3 Protein Subtomogram Dataset
 We created a realistic simulated cryo-ET subtomogram dataset of 18,000 subtomograms of size $32^{3}$.
@@
 Moreover, cellular cryo-ET is the primary method to capture these proteins inside the cells in their native state [51].
+]Refer to caption
+Figure 13: 3D slice-by-slice visualization of Nucleosome subtomograms. Each subtomogram is associated with the PDB ID of the original structure.
+Figure 14: Isosurface visualization of Nucleosome Density Maps. Each density map slightly varies in terms of conformation. They are associated with the PDB IDs in the figure.
+Figure 15: 3D slice-by-slice visualization of Spike Protein Subtomograms. Each subtomogram is associated with the PDB ID of the original structure.
+Figure 16: 3D slice-by-slice visualization of FAS subtomograms. Each subtomogram is associated with the PDB ID of the original structure.
 For each protein class, we collected 6 different protein structures from the RCSB PDB website [30].
@@
 For nucleosomes, we collected PDB IDs ‘2pyo’, ‘7kbe’, ‘7pex’, ‘7pey’, ‘7xzy’, and ‘7y00’.
-All of these are different conformations of nucleosomes that slightly vary in composition (Figure LABEL:fig:supp:nuc-dmaps).
+All of these are different conformations of nucleosomes that slightly vary in composition (Figure 14).
 For sars-cov-2 spike proteins, we collected PDB IDs ‘6vxx’, ‘6vyb’, ‘6xr8’, ‘6xra’, ‘6zox’, and ‘6zp0’.
@@
 We uploaded the entire dataset anonymously at https://zenodo.org/records/11244440 under CC-BY-SA license.
-3D slice-by-slice visualizations for sample subtomograms for nucleosomes, spike proteins, and FAS units are provided in Figure 6.
-We trained DualContrast and baseline approaches against the protein subtomogram dataset.
+Sample subtomograms for nucleosomes, spike proteins, and FAS units are provided in Figure 13, Figure 15, and Figure 16 respectively.
+The figures show 3D slice-by-slice visualization for each conformation of the corresponding protein.
 We could not train VITAE [1] on subtomogram datasets since it did not define any transformation for 3D data.
 Designing CPAB transformation for 3D data by ourselves was challenging.
-Among the other baseline approaches, spatialVAE could not distinguish the protein identities with high heterogeneity at all, which is evident by its embedding space UMAP (Figure 7).
-Only Harmony provided somewhat plausible results apart from DualContrast.
-Figure 7: UMAP of content factor learned by SpatialVAE [2] on our subtomogram dataset. The protein identities are indistinguishable from the embedding space.
-We observed that the DualContrast content factor perfectly clustered the three protein identities from the dataset (GMM clustering accuracy $0.99$), profoundly disentangling the transformations (Figure 8(b)).
-We do not see such perfect clustering for Harmony (Figure 8(a)); rather, protein identities with very different compositions get mixed up due to their conformational variation (GMM clustering accuracy $0.70$).
-We further observe that DualContrast transformation embedding clusters have similar nucleosome conformations, showing disentanglement of conformation from protein identity.
-On the other hand, in Harmony, the transformation factor only represents rotation and translation by design and can not capture nucleosome conformations.
-Moreover, even its content embedding for manually selected (not through automatic clustering) nucleosome subtomograms can not show good clustering of conformations (Figure 8(b)).
-Figure 8: Disentanglement of composition and conformations in cellular subtomogram dataset.
-(a) UMAP embedding of the content factor of all subtomograms (top) and only nucleosome subtomograms (bottom) with Harmony [4] (b) UMAP embedding of the content factor (top) and transformation factor (bottom) obtained by DualContrast.
-Ablation Study: To evaluate the individual contribution of the contrastive losses, we conduct both quantitative and qualitative ablation analyses of DualContrast.
+However, we trained SpatialVAE and Harmony as baselines against our subtomogram dataset.
+Between these two, spatialVAE could not distinguish the protein identities with high heterogeneity at all, which is evident by its embedding space UMAP (Figure 5).
+Only Harmony and DualContrast showed plausible result, where DualContrast showing much superior disentanglement than Harmony (Figure 5).
+A.4.4 Ablation Study
+To evaluate the individual contribution of the contrastive losses, we conduct both quantitative and qualitative ablation analyses of DualContrast.
 We trained (1) DualContrast without any contrastive loss, which is basically a VAE with two latent spaces, (2) DualContrast with only $L=L_{\text{VAE}}+L_{\text{con (z)}}$, and (3) $L=L_{\text{VAE}}+L_{\text{con (c)}}$.
 We qualitatively and quantitatively evaluated each model.
-We show the quantitative results in Table 1 and qualitative results of MNIST in Figure 9.
+For (1), the $D_{\text{score}}$ is almost the same for both c and z codes, indicating equal predictivity of the digit classes by both codes.
+This is obvious given that the model has no inductive bias to make different codes capture different information.
+In model (2), using contrastive loss w.r.t.
+only z factor makes it uninformative of the data.
+Thus, it provides a small $D(c|z)$ score as desired, but the changing z does not affect the image generation (Fig.
+7).
+On the other hand, in the model (3), using contrastive loss w.r.t.
+only c gives a less informative c factor, a lower $D(c|c)$ score, and a higher $D(c|z)$ score, contrary to what is desired.
+These results indicate that contrastive loss w.r.t to both codes is crucial for the desired disentanglement.
+Furthermore, we investigated whether using only positive pairs or negative pairs for both codes is sufficient for disentanglement.
+Nevertheless, we found that both leads to suboptimal disentanglement.
+If only negative pairs are used, only rotation is disentangled.
+If only positive pairs are used, then the transformation code becomes uninformative of the data, similar to the degenerate solution.
+We provide further quantitative results on the ablation study in Figure 17.
+The image grids show decoder-generated images where the content factor is used from the corresponding topmost row, and the transformation factor is used from the corresponding leftmost column image.
+((a))
+((b))
+((c))
+((d))
 Figure 17: Content-transformation transfer results from ablation analysis.
-(a) and (b) shows the results when the model is trained with $L_{\text{VAE}}+L_{\text{con (z)}}$ and $L_{\text{VAE}}+L_{\text{con (c)}}$ respectively.
-For (1), the $D_{\text{score}}$ is almost the same for both content and transformation factors, indicating equal predictivity of the digit classes by both factors.
-This is obvious given that the model has no inductive bias to make different factors capture different information.
-In model (2), using contrastive loss w.r.t.
-only transformation factor $z$ makes it uninformative of the data.
-Thus, it provides a small $D(c|z)$ score as desired, but the changing $z$ does not affect the image generation (Figure 9).
-On the other hand, in the model (3), using contrastive loss w.r.t.
-only content factor $c$ gives a less informative content factor, a lower $D(c|c)$ score, and a higher $D(c|z)$ score, contrary to what is desired.
-These results indicate contrastive loss w.r.t to both factors is crucial to get the desired disentanglement.
-5 Limitation
-Similar to existing content-transformation disentanglement methods [2, 4, 1, 8, 12], DualContrast is applicable to data samples with single objects.
-This is reasonable, given content and transformation can not be defined for data with multiple objects.
-Moreover, we limited our study only to shape-focused image datasets, motivated by the problem of conformation disentanglement of protein subtomograms.
-Consequently, we only used rotation to create contrastive pairs, which worked pretty well in our cases.
-For image datasets where appearance or illumination plays a crucial role, using only rotation to create contrastive pairs may not work well.
-However, disentanglement in natural image datasets is a different scenario that is outside of the scope of this paper.
-6 Conclusion
-This work focuses on a challenging setting of the unsupervised content-transformation disentanglement problem in shape-focused image datasets, where the transformation latent factor is not explicitly parameterized.
-To tackle this problem, we propose a novel method termed DualContrast.
-DualContrast employs generative modeling with a novel contrastive learning strategy that creates positive and negative pairs for content and transformation latent factors.
-Our extensive experiments show DualContrast’s effective disentanglement of challenging transformation factors across various shape-based image datasets; including simulated cellular subtomograms, where it solved the unexplored problem of isolation of protein conformations from protein identities, as proof of principle.
-7 Acknowledgement
-This work was supported in part by U.S.
-NIH grants R01GM134020 and P41GM103712, NSF grants DBI-1949629, DBI-2238093, IIS-2007595, IIS-2211597, and MCB-2205148.
-This work was supported in part by Oracle Cloud credits and related resources provided by Oracle for Research, and the computational resources support from AMD HPC Fund.
-MRU were supported in part by a fellowship from CMU CMLH.
+(a), (b), (c), and (d) show the results when the model is trained with $L_{\text{VAE}}$, $L_{\text{VAE}}+L_{\text{con (z)}}$, $L_{\text{VAE}}+L_{\text{con (c)}}$ and full objective respectively.